### v1

# Active Preference Learning for Ordering Items In- and Out-of-sample

Abstract
Learning an ordering of items based on noisy pairwise comparisons is useful when item-specific labels are difficult to assign, for example, when annotators have to make subjective assessments. Algorithms have been proposed for actively sampling comparisons of items to minimize the number of annotations necessary for learning an accurate ordering. However, many ignore shared structure between items, treating them as unrelated, limiting sample efficiency and precluding generalization to new items. In this work, we study active learning with pairwise preference feedback for ordering items with contextual attributes, both in- and out-of-sample. We give an upper bound on the expected ordering error incurred by active learning strategies under a logistic preference model, in terms of the aleatoric and epistemic uncertainty in comparisons, and propose two algorithms designed to greedily minimize this bound. We evaluate these algorithms in two realistic image ordering tasks, including one with comparisons made by human annotators, and demonstrate superior sample efficiency compared to non-contextual ranking approaches and active preference learning baselines.
1 Introduction
The success of supervised learning is built on annotating items at great volumes with small error. For subjective assessments, however, assigning a value from an arbitrary rating scale can be difficult, causing many to favor preference feedback from pairwise comparisons  (Phelps et al., 2015; Yannakakis and Martínez, 2015; Christiano et al., 2017; Zhu et al., 2023). Preference feedback is sufficient to learn an ordering of items (Fürnkranz and Hüllermeier, 2003), but for $n$ items, there are $\mathcal{O}(n^{2})$ possible pairs of items to compare. A common solution is to use crowd-sourcing  (Chen et al., 2013; Yang et al., 2021; Larkin et al., 2022), but what if the task requires certain expertise and annotations are expensive to collect? This is the case in the field of medical imaging, where annotations require trained radiologists (Phelps et al., 2015; Jang et al., 2022; Lidén et al., 2024, Appendix A). To be cost-efficient, we must select comparisons carefully, to minimize the annotations needed to learn a good ordering.
There is a large literature on ordering a set of items $\mathcal{I}$ using only noisy pairwise comparisons of actively chosen item pairs (Herbrich et al., 2006; Maystre and Grossglauser, 2017; Heckel et al., 2018; Lou et al., 2022). While robust against noise and more efficient than exhaustive comparison, these algorithms ignore any context attributes that the items may have. Say that a radiologist wants to quantify the expression of a disease in a collection of X-ray images $\mathcal{I}$. Ranking these images according to their severity $y_{i}$ using purely preference-based algorithms ignores the content $x_{i}$ of the X-rays, sampling images $i,j\in\mathcal{I}$ for comparison based only on the current estimated severity and previous feedback. However, the content of the images reveal similarities between X-rays which can inform an ordering strategy.
In applications, the set $\mathcal{I}$ we want to order is often larger than the set of items we observe during training—we may want to rank new images in relation to previous ones. However, methods based only on preference feedback cannot generalize what they learned to new items. As an alternative, active learning algorithms for classification can be used for ordering by fitting a map from pairs of contexts $x_{i},x_{j}$ to the comparison $i\succ_{?}j$, that can be applied to new items (Houlsby et al., 2011; Kirsch et al., 2019; Bengs et al., 2022). However, to the best of our knowledge, there is no analysis that justifies using these algorithms to learn a complete ordering or gives guarantees for their performance.
Contributions.
In this work, we study active preference learning algorithms for ordering sets of items based on noisy pairwise comparisons. We propose the use of a contextual logistic preference model to support efficient ordering in-sample as well as generalization to new items. Our analysis yields the first bound on the expected ordering error achievable under this model, provided a fixed budget of comparisons (Section 4). This result justifies an active sampling principle based on the epistemic and aleatoric uncertainty in the outcome of a comparison between two items. We use this in one greedy, deterministic, and one Bayesian, stochastic algorithm (Section 5) applied to both a fully contextual preference model, and a hybrid variant with per-item parameters, aimed at overcoming potential model misspecification (Section 5.3). We evaluate these models and sampling strategies in two realistic image ordering tasks, one where the preference feedback is simulated using a logistic model, and one where the comparisons have been performed by human annotators (Section 6). We show that our sampling strategy compares favorably to active preference learning and bandit baselines, and that our hybrid model benefits both from the low variance of contextual models and the low bias of fitting per-item parameters. This manifests in faster convergence in-sample, better generalization to new items, and more efficient tuning when continuing learning on new items.
2 Ordering with active preference learning
Our goal is to learn an ordering of items $\mathcal{I}$ according to an unobserved score $y_{i}\in\mathbb{R}$, defined for each item $i\in\mathcal{I}$. The ground-truth ordering of $\mathcal{I}$ is determined by a comparison function $\pi_{ij}\coloneqq\mathds{1}[y_{i}>y_{j}]$ such that $\pi_{ij}=1$ indicates that $i$ is ranked higher than $j$. We assume there are no ties.
We define the ordering error $R_{\mathcal{I}}(h)$ of a learned comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$ as the frequency of pairwise inversions under a uniform distribution of item pairs,
$R_{\mathcal{I}}(h)=\frac{2}{n(n-1)}\sum_{i\neq j\in\mathcal{I}}\mathds{1}[h(i,j)\neq\pi_{ij}]~{},$ (1)
where $n=|\mathcal{I}|$. This error is equivalent to the normalized Kendall’s Tau distance (Kendall, 1948). Our analysis may be readily generalized to non-uniform distributions.
We learn $h$ from preference feedback—noisy pairwise comparisons $C_{ij}\in\{0,1\}$ for items $(i,j)$ related to their score, for example, provided by one or more human annotators. $C_{ij}=1$ indicates that the annotator perceived that item $i$ has a higher score than $j$, i.e., that they prefer $i$ over $j$. Our goal is to minimize the ordering error $R_{\mathcal{I}}(h)$ for a fixed budget $T\geq 1$ of adaptively chosen comparisons.
We are particularly interested in algorithms that are guaranteed to generalize out-of-sample to unseen items. In this setting, only a subset of $\mathcal{I}_{D}\subseteq\mathcal{I}$ is available during learning, and the learner must generalize to $\mathcal{I}\setminus\mathcal{I}_{D}$. Ordering algorithms based only on preference feedback cannot solve this problem since observed comparisons are uninformative of new items. To overcome this, we let each item $i\in\mathcal{I}$ be endowed with item-specific attributes $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. The active preference learning scenario proceeds as follows:
  1. 1.
    A learner is given an annotation budget $T$, a pool of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and per-item attributes $x_{i}$ for $i\in\mathcal{I}_{D}$.
  2. 2.
    Over rounds $t=1,...,T$, the learner requests a comparison of two items $i_{t},j_{t}\in\mathcal{I}_{D}$ according to a sampling criterion and receives noisy binary preference feedback $c_{t}\sim p(C_{ij})$, independently of previous comparisons.
  3. 3.
    After $T$ rounds, the learner returns an estimated comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$.
We denote the history of accumulated observations (pairs of items and comparisons) up until and including time $t$ by $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$. To learn a good model by time $T$, an active learner must use all information in $D_{t}$ to inform its sampling criterion. A natural baseline is the uniform criterion, which selects any pair $i\neq~{}j\in\mathcal{I}_{D}$ with equal probability, ignorant of $D_{t}$. Uniform sampling gives equal weight to item pairs with similar scores and pairs with vastly different scores, which may result in many comparisons having minimal impact on the final estimate $h$.
Logistic preference feedback. We analyze the case where comparisons $C_{ij}$ are generated according to a logistic model applied to the difference between item scores, $p(C_{ij}=1)=\sigma(y_{i}-y_{j})~{},$ where $\sigma(z)=1/(1+e^{-z})$ is the logistic function and $\dot{\sigma}(z)$ its derivative at $z$. Note that $\dot{\sigma}(z)$ is also the variance of a Bernoulli variable parameterized by $\sigma(z)$.
We are particularly interested in the case where $y_{i}$ is a linear function of item attributes, $y_{i}=\theta_{*}^{\top}x_{i}~{},$ with $\theta_{*}\in\mathbb{R}^{d}$ the ground-truth coefficients. In this case, comparisons are determined by a logistic regression model applied to the difference vector between item attributes $z_{ij}\coloneqq x_{i}-x_{j}$,
$p(C_{ij}=1)=\sigma(\theta_{*}^{\top}z_{ij})~{}.$ (2)
This model is a special case of the family of parametric models considered in Heckel et al. (2018), and is equivalent to the logistic model often used in learning to rank (Burges et al., 2005). It satisfies linear (and thus strong) stochastic transitivity (Oliveira et al., 2018). A similar score function is assumed also in Chu and Ghahramani (2005).
3 Related work
Preference learning (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005) is related to the problem of learning to rank (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012). Solving preference learning using as few, adaptively chosen comparisons as possible may be posed as an active learning (Brinker, 2004; Long et al., 2010; Silva et al., 2014; Ling et al., 2020) or bandit problem. Non-contextual active learners, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018), and Probe-Rank (Lou et al., 2022) produce in-sample orderings, but must be updated if new items are to be ranked. This is mitigated by contextual algorithms that exploit structure between items, such as BALD (Houlsby et al., 2011). BALD, when applied to preference learning, samples items by maximizing the expected information gain, a commonly used criterion in Bayesian optimal design (Lindley, 1956). Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information. Chaudhuri et al. (2015) analyze active learning for general maximum likelihood estimation, and Ash et al. (2021) extend this idea for models parameterized by deep neural networks.
Bandit algorithms with relative or dueling feedback (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022), have been proposed both in contextual (Dudík et al., 2015) and non-contextual settings (Yue et al., 2012). Bengs et al. (2022) proposed CoLSTIM, a contextual dueling bandit for regret minimization under linear stochastic transitivity, matching the model in (2). Di et al. (2023) gave variance-aware regret bounds for this setting. However, as argued in Appendix E, algorithms that identify the top-$k$ items as fast as possible, such as pure exploration bandits (Fang, 2022; Jun et al., 2021), can be arbitrarily bad at learning a full ordering. Related are also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits, and Wu et al. (2023b) who minimize the Borda regret. Our analysis uses techniques from generalized linear bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020). For an extended review of related work, see Appendix B.
4 A bound on the ordering error
We give an upper bound on the ordering error $R_{\mathcal{I}}(h)$ for a hypothesis $h$ fit using a fixed budget of comparisons, $T$. We restrict our analysis to the logistic model in (2) and denote by $R(\theta)\equiv R_{\mathcal{I}}(h_{\theta})$ the risk of the hypothesis defined by $h_{\theta}(i,j)=\mathds{1}[\theta^{\top}z_{ij}>0]$. Let $\theta_{t}$ be the maximum-likelihood estimate (MLE) of the true parameter $\theta_{*}$, defined in (9) in the Appendix, fit to $t$ rounds of feedback, $D_{t}$.
Fix $i,j\in\mathcal{I}$ and recall that $z_{ij}=x_{i}-x_{j}$. For convenience, we define $z_{t}\equiv z_{i_{t}j_{t}}$ as the difference between attributes for the pair of items selected at round $t$. Now, let $\Delta_{ij}>0$ be a lower bound on the margin of comparison such that $\left|\sigma(z_{ij}^{\top}\theta_{*})-1/2\right|>\Delta_{ij}$ for all $i,j$. Further, define $\Delta_{*}=\min_{ij}\Delta_{ij}$. Let $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$ be the Hessian of the negative log-likelihood of observations at time $t$ under (2), given the parameter $\theta$. We define
$\tilde{\mathbf{H}}_{t}(\theta):=\sum_{i,j}w_{t}(i,j)\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}$
with $w_{t}(i,j))=N_{t}(i,j)/t$ where $N_{t}(i,j)$ denotes the number of times the pair $(i,j)$ has been queried up to time $t$. Hence, $\mathbf{H}_{t}(\theta_{t})=t\tilde{\mathbf{H}}_{t}(\theta_{t})$. We further note that $\mathbf{H}_{t}(\theta)$ is the observed Fisher information, and define the norm $||x||_{V}$, for a square matrix $V$, as $\|x\|_{V}=\sqrt{x^{\top}Vx}$.
We make the following assumptions for our analysis:
Assumption 1.
$\theta_{*}$ satisfies $\|\theta_{*}\|_{2}\leq S$ for some $S>0$.
Assumption 2.
$\forall i\in\mathcal{I}$, we have $\|x_{i}\|_{2}\leq Q$ for $Q>0$.
Assumption 3.
$\mathbf{H}_{T}(\theta_{T})$ and $\mathbf{H}_{T}(\theta_{*})$ have full rank and their minimum eigenvalues are larger than some $\lambda_{0}>0$.
Assumptions 1 and 2 are required for our analysis but the learner does not have to know these bounds. Assumption 3 is naturally satisfied, for sufficiently large $T$, by most reasonable sampling strategies, or can be ensured by allowing for a burn-in phase of $O(d)$ uniformly drawn samples in the beginning of an adaptive strategy.
We start by stating the following concentration result for the deviation of $\sigma(z_{ij}^{\top}\theta_{T})$ from the true probability $\sigma(z_{ij}^{\top}\theta_{*})$. The proof of Lemma 1 is found in Appendix D and builds on results for optimistic algorithms in logistic multi-armed bandits (Filippi et al., 2010; Faury et al., 2020).
Lemma 1 (Concentration Lemma).
Define, for all pairs of items $i,j\in\mathcal{I}$, and any $\Delta>0$,
$\alpha_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta^{2}T}{8dC_{1}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T}))}\right)^{2}}\Bigg{)},\;\;\beta_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta T}{dC_{1}\left(\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)^{2}}\Bigg{)}~{}.$
Then, if $\alpha\coloneqq\alpha_{ij}(\Delta),\beta\coloneqq\beta_{ij}(\Delta)$ and $\alpha,\beta\leq\frac{1}{4dT}$,
$P\left(|\sigma(z_{ij}^{\top}\theta_{*})-\sigma(z_{ij}^{\top}\theta_{T})|>\Delta\right)\leq 2dT\left(\alpha+\beta\right).$
Here $C_{1}$ is a constant depending on $S,\lambda_{0},Q$ from Assumptions 13 (see Appendix D for definition).
The concentration result in Lemma 1 is verifiable (given by observables) since the upper bound depends only on the maximum likelihood estimate $\theta_{T}$ at time $T$, not on $\theta_{*}$. We present a sharper, unverifiable bound in Appendix D which instead depends on $\theta_{*}$ but does not suffer from the explicit scaling with $d$ in the definitions of $\alpha$ and $\beta$. The bound in Lemma 1 can also be expressed in terms of $\mathbf{H}^{-1}_{T}(\theta_{T})$ by using the equality $||z_{ij}||^{2}_{\mathbf{H}^{-1}_{T}(\theta_{T})}=\frac{1}{T}||z_{ij}||^{2}_{\hat{\mathbf{H}}^{-1}_{T}(\theta_{T})}$
As long as our sampling strategy ensures that the minimum eigenvalue of $\tilde{\mathbf{H}}_{t}(\theta_{t})$ does not tend to zero, i.e., the strategy is strongly consistent (Chen et al., 1999), we have $\alpha_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}^{2}T/(\dot{\sigma}(z_{ij}^{\top}\theta_{T})^{2}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})\right)$ and $\beta_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}T/\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)$. Since $\Delta^{2}_{ij}<\Delta_{ij}<1/2$ by definition, we can view $\alpha$ as the first-order term and $\beta$ as the second-order term of our bound. Lemma 1 formally captures the intuition that it should be easier to sort when annotations contain little noise, i.e., $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$ is small. Especially, we observe $\dot{\sigma}(z_{ij}^{\top}\theta_{T})\approx 0$ for pairs where $\Delta_{ij}$ is sufficiently large, causing the first-order term to vanish, leaving us with the faster decaying second-order term $\beta$.
Lemma 1 also tells us that the hardest pairs to guarantee a correct ordering for are the ones with both high aleatoric uncertainty under the MLE model, e.g., where annotators disagree or labels are noisy, captured by $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$, as well as high epistemic uncertainty captured by $||z_{ij}||_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$.
A direct consequence of Lemma 1 is the following bound,
$\mathbb{E}[R(\theta_{T})]\leq\sum_{i\neq j}\frac{2\min\left\{2dT\left(\alpha_{ij}(\Delta_{ij})+\beta_{ij}(\Delta_{ij})\right),1\right\}}{n(n-1)}.$
The RHS in the above inequality can be bounded further by utilizing that $\Delta_{ij}\geq|i-j|\Delta_{*}$. Together with Markov’s inequality, this yields the following bound on $P\left(R(\theta_{T})\geq\epsilon\right)$.
Theorem 1 (Upper bound on the ordering error).
Let $\alpha_{*}:=\max_{i\neq j}\alpha_{ij}(\Delta_{*})$ and $\beta_{*}:=\max_{i\neq j}\beta_{ij}(\Delta_{*})$, with $\alpha,\beta$ from Lemma 1. Then, for $\alpha_{*},\beta_{*}\leq\frac{1}{4dT}$ and any $\epsilon\in(0,1)$, the ordering error $R(\theta_{T})$ satisfies
$P(R(\theta_{T})\geq\epsilon)$ $\leq\frac{4dT}{\epsilon n}\left(\left(\alpha_{*}^{-1}-1\right)^{-1}+\left(\beta_{*}^{-1}-1\right)^{-1}\right)\approx\frac{4dT}{\epsilon n}\left(\alpha_{*}+\beta_{*}\right)~{},$
where $\alpha_{*}$ and $\beta_{*}$ decay exponentially with $T$.
A proof is given in Appendix D.
Theorem 1 suggests that the probability of $R(\theta_{T})\geq\epsilon$ decays exponentially with a rate that depends on the quantities $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$ and $\max_{i,j}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$. Both quantities are random variables that depend on the particular sampling strategy that yields $\mathbf{H}_{t}$. Focusing on the leading term, $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$, Theorem 1 suggests that an active learner should gather data to minimize this quantity and obtain the smallest possible bound.
The factor $\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$, is related to active sampling criteria based on Fisher information (Kirsch and Gal, 2022); it is the weighted norm of $z_{ij}$ w.r.t. the inverse of the observed Fisher information. It controls the shape of the confidence ellipsoid around $\theta_{T}$ and the width of the confidence interval around $\theta_{T}^{\top}z_{ij}$. The leading term in Theorem 1 re-scales this quantity with aleatoric noise under the MLE estimate $\theta_{T}$.
Regularized maximum-likelihood.
Throughout this section, we have assumed that $\theta_{T}$ is the maximum likelihood estimate and that $\mathbf{H}(\theta_{T})$ has full rank. This can be relaxed by considering $\ell_{2}$ (Ridge) regularization where $\theta_{\lambda_{0},T}$ is the optimum of the regularized log-likelihood with regularization $\lambda_{0}\mathbf{I}$ and $\mathbf{H}_{\lambda_{0}}(\theta_{\lambda_{0},T})=\sum_{s=1}^{T}\dot{\sigma}(z_{s}^{\top}\theta_{\lambda_{0},T})z_{s}z_{s}^{\top}+\lambda_{0}\mathbf{I}$. The same machinery used to prove Lemma 1 (Filippi et al., 2010; Faury et al., 2020) can be applied to this regularized version and result in small changes to the final bound.
Other link functions.
We have focused on the logistic regression model as it is often used to model preferences. It is possible to derive similar results for other generalized linear models, with other link functions, $\mu(z_{ij}^{\top}\theta_{*})$. For this, we could use the inequality $\mathbf{H}(\theta)\geq\kappa^{-1}\mathbf{V}$ with $\mathbf{V}=\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ and $\kappa\geq 1/\min_{z_{ij}}\dot{\mu}(z_{ij}^{\top}\theta_{*})$. To the best of our knowledge, this would yield a scaling of $\sim\exp(-\Delta^{2}T/\kappa)$ where, unfortunately, $\kappa$ might be very large. For discussion on the dependence on $\kappa$ in generalized linear bandits, see Lattimore and Szepesvári (2020, Chapter 19).
5 Algorithms
We present two active preference learning algorithms for ordering a list of items: GURO, a deterministic algorithm based on greedy minimization of the bound in Theorem 1, and BayesGURO, a stochastic version where we sample $\theta$ from a posterior distribution. Both are presented in Algorithm 1. As we will see, they both lead to similar sampling criteria. We begin by studying preference models in the form $\sigma(\theta^{\top}z_{ij})$ and return in Section 5.3 to variations on the model parameterization.
Algorithm 1 Greedy Uncertainty Reduction for Ordering
1:Training items $\mathcal{I}_{D}$, attributes ${\bf X}=\{x_{i}\}_{i\in\mathcal{I}_{d}}$
2:Initialize $\theta_{0}$
3:for $t=1,...,T$ do
4:     Draw $(i_{t},j_{t})$ according to (4) or (5)
5:     Observe $c_{t}$ from noisy comparison (annotator)
6:     $D_{t}=D_{t-1}\cup\{i_{t},j_{t},c_{t})\}$
7:     if GURO then
8:          Let $\theta_{t}=\mathrm{MLE}(D_{t})$,    (9) in the Appendix
9:     else if BayesGURO then
10:          Let $\theta_{t}=\mathrm{MAP}(D_{t})$,    (10) in the Appendix
11:     end if
12:end for
13:Return $h_{T}$
5.1 Greedy worst-pair uncertainty reduction
The main component of the bound in Theorem 1 to be controlled by an active learning algorithm is
$\max_{i,j\in\mathcal{I}}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\mathbf{H}^{-1}_{T}(\theta_{T})}~{},$ (3)
which represents the highest uncertainty in the comparison of any items $i,j\in\mathcal{I}$ under the model $\theta_{T}$. A smaller value of (3) yields a smaller bound and a stronger guarantee.
Recall that, for any $t=1,...,T$, $\theta_{t}$ is the MLE estimate of the ground-truth parameter $\theta_{*}$ with respect to the observed history $D_{t}$. Both factors in (3) are determined by the sampling strategy that yielded the item pairs $(i_{t},j_{t})$ in $D_{T}$ and, therefore, $\mathbf{H}_{T}$ and $\theta_{T}$ (the results of comparisons $c_{ij}$ are outside the control of the algorithm, but $z_{ij}$ are known).
Direct minimization of (3) is not feasible without access to comparisons $c_{ij}$ and their likelihood under $\theta_{T}$. Instead, we adopt a greedy, alternating approach: In each round, a) a single pair is sampled for comparison by maximizing (3) under the current model estimate, and b) $\theta_{t}$ is recomputed based on $D_{t}$. Specifically, at $t=1,...,T$, we sample,
$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i\neq j}\dot{\sigma}(z_{ij}^{\top}\theta_{t-1})\|z_{ij}\|_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}~{}.$ (4)
We refer to this criterion as GURO for Greedy Uncertainty Reduction for Ordering, since sampling in this way reduces the uncertainty of $\theta_{t}$ in the direction of $z_{ij}$. To see this, consider the change of $\mathbf{H}_{t}(\theta_{t})$ after a single play of $i_{t},j_{t}$. The Sherman-Morrison formula (Sherman and Morrison, 1950) yields, with $z_{t}\coloneqq z_{i_{t}j_{t}}$,
$\mathbf{H}_{t}^{-1}(\theta_{t-1})$ $=\mathbf{H}_{t-1}^{-1}(\theta_{t-1})-\underbrace{\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\frac{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})z_{t}z_{t}^{\top}\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}{1+\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\|z_{t}\|^{2}_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}}}_{=:{\color[rgb]{1,0,0}\xi}}$
and thus, for all $i<j\in\mathcal{I}$, with $\mathbf{H}_{t-1}=\mathbf{H}_{t-1}(\theta_{t-1})$, it holds that $||z_{ij}||^{2}_{\mathbf{H}_{t}^{-1}(\theta_{t-1})}=||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}-||z_{ij}||^{2}_{{\color[rgb]{1,0,0}\xi}}\leq||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}~{}.$ The inequality is strict for the pair $i_{t},j_{t}$ in (4). As $\theta_{t}$ converges to $\theta_{*}$, this pair becomes representative of the maximizer of (3) provided there is no major systematic discrepancy between $\mathcal{I}_{D}$ and $\mathcal{I}$.
A drawback of a deterministic greedy criterion like (3) is that it cannot readily be applied in batched settings since it will sample the same pair of items throughout the batch. Next, we offer a Bayesian stochastic alternative.
5.2 Stochastic Bayesian uncertainty reduction
Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$. In our second algorithm, we sample items $i_{t},j_{t}$ based on a stochastic estimate of the posterior variance of predictions for their comparison,
$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i<j}\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]~{}.$ (5)
Here, $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ is an estimate of the variance of probabilities $\sigma(\theta^{T}z_{ij})$, computed from finite samples drawn from the posterior of $\theta$. Estimating the variance in this way both i) allows for tractable implementation, and ii) induces controlled stochasticity in the selection of item pairs. We refer to the sampling criterion in (5) as BayesGURO.
For the logistic model considered in this section, using Laplace approximation with a Normal prior $\mathcal{N}(0,\mathbf{H}^{-1}_{B,0})$ on $\theta$, the Bayesian criterion in (5) is related to the criterion in (4) through the first-order Taylor expansion of the variance:
$\mathbb{V}_{\theta\mid D_{t}}(\sigma(\theta^{\top}z_{ij}))$ $\approx(\dot{\sigma}(\mathbb{E}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]))^{2}\mathbb{V}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]$
$=(\dot{\sigma}(\theta_{B,t}^{\top}z_{ij})||z_{ij}||_{\mathbf{H}^{-1}_{B,t}(\theta_{B,t})})^{2}~{},$
where $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ and $\mathbf{H}_{B,t}$ is the Hessian adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$ (further described in Appendix C.2). Thus, to a first-order approximation, for a large number of posterior samples, the two active learning criteria are equivalent, save for the influence of the prior. In practice, we find that the Bayesian variant lends itself well to sequential updates of the posterior. The choice of prior $p(\theta)$, which could be useful under strong domain knowledge, and the stochasticity of using few posterior samples to approximate $\mathbb{V}$ make the two criteria distinct.
5.3 Preference models
We consider logistic preference models aimed at mitigating different sources of error. In all variants, we let $h(i,j)=\mathds{1}[f(i,j)>0]$ for some function $f$ approximating the difference in score between items $i$ and $j$. Throughout, we estimate $f$ using probabilistic modeling, with a logistic likelihood function $\sigma(f(i,j))\approx p(C_{ij}=1)$, matching (2).
Our default model (see (2) and Section 4) is fully contextual,
$f_{\theta}(x_{i},x_{j})=\theta^{\top}(x_{i}-x_{j})~{}.$ (6)
A strength of this model is that the variance in its estimates grows with $d$, but not with $n=|\mathcal{I}|$. As we see empirically (e.g., in Figure 4), this results in much quicker convergence than non-contextual methods. This fully contextual model also generalizes to unseen items in $\mathcal{I}$ under modest assumptions on how $\mathcal{I}_{D}$ was sampled.
The model’s main drawback is two other sources of error: model misspecification bias (error due to the chosen functional form), and noise (error due to insufficiency of $X$ in predicting $C$). The former can be mitigated by applying the linear model to a representation function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$,
$f_{\theta}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j}))~{}.$
A good representation $\phi$, e.g., from a foundation model, can mitigate misspecification bias, and enable using different data modalities of input. As demonstrated in Figure 7 in the Appendix, even a representation pre-trained for a different task can perform much better than a random initialization. [^1] It is also feasible to update representations during exploration (Xu et al., 2022), but we do not consider that here.
Noise due to insufficiencies in $X$ cannot be mitigated by a representation $\phi(x)$; If $X$ does not sufficiently cover aspects considered by annotators, even the Bayes classifier $\mathds{1}[p(C_{ij}\mid X_{i},X_{j})>0.5]$ will have a large ordering error. However, for in-sample ordering, we can introduce per-item parameters $\zeta_{i}\in\mathbb{R}$ to the score function, one for each item $i\in\mathcal{I}_{D}$, that compensate for both model misspecification and for uninformative attributes,
$f_{\theta,\boldsymbol{\zeta}}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j}))+(\zeta_{i}-\zeta_{j})~{}.$ (7)
Here, $\zeta_{i}-\zeta_{j}$ corrects the residual of the fully contextual model, which can be assumed small if a) the context captures the most relevant information about the ordering, and b) the functional form $\theta^{\top}\phi(x_{i})$ is nearly well-specified. $\zeta_{i}-\zeta_{j}$ alone are sufficient to learn a perfect in-sample ordering, at the cost of increased variance (the full dimension is now $n$ instead of $d$) and poor generalization (since $\zeta_{i}$ are unknown for items $i\not\in\mathcal{I}_{D}$). In experiments, the hybrid model in (7) exhibits the best of both contextual and per-item parameterizations: When $\phi$ is poor, the model recovers and competes with non-contextual models (Figure 7); when $\phi$ is good, convergence is as quick as fully contextual models.
6 Experiments
We evaluate the GURO variants in two image ordering tasks, one with logistic (synthetic) preference feedback, and one based on real-world feedback from human annotators. We provide a synthetic experiment in Appendix F.3 that includes empirical estimates of the bound in Theorem 1.
Baselines
We select diverse baseline algorithms for active sampling of comparisons: BALD (Houlsby et al., 2011), CoLSTIM (Bengs et al., 2022), TrueSkill (Graepel, 2012; Hees et al., 2016), Uniform sampling, and a version of GURO that ignores aleatoric uncertainty (referred to as NormMin).
Bayesian Active Learning by Disagreements (BALD) is, a priori, considered the strongest baseline since it is a contextual active learning algorithm, unlike the others. When applied to preference learning, BALD determines the next comparison by maximizing the following criterion
$\operatorname*{arg\,max}_{i\neq j\in\mathcal{I}_{D}}H[C_{ij}\mid z_{ij},D_{t}]-\underset{p(\theta\mid D_{t})}{\mathbb{E}}[H[C_{ij}\mid z_{ij},\theta_{t}]]~{},$ (8)
where $H$ is the conditional entropy. The criterion aims to greedily maximize the decrease in posterior entropy, which amounts to reducing the epistemic uncertainty. The second term in (8) will downplay the influence of the aleatoric uncertainty. This is not always beneficial, as suggested by our results in Section 4, since learners may require several comparisons of high-uncertainty pairs to get the order right.
CoLSTIM was developed for regret minimization and is not expected to perform well here, but is included to illustrate the mismatch between regret minimization and our setting. Uniform sampling is oblivious to the observed data and represents a non-active learner, but remains a competitive baseline in active learning tasks (Yang and Loog, 2018).
TrueSkill (Herbrich et al., 2006) is a non-contextual skill-rating system that models each item as a Gaussian distribution without leveraging item attributes. We use the sampling rule from Hees et al. (2016), designed for ordering. Finally, to illustrate the importance of accounting for aleatoric uncertainty, we use a version of GURO called NormMin that ignores the $\dot{\sigma}(z_{ij}^{\top}\theta_{t})$ term and plays the pair maximizing $\|z_{ij}\|_{\mathbf{H}_{t}^{-1}(\theta_{t})}$, i.e., it minimizes the second-order term in Lemma 1. See Appendix F.3 for a detailed comparison.
(a) Kendall’s tau distance between estimated and true orderings.
(b) Mean generalization error with the 95% confidence interval.
Figure 1: X-RayAge. Performance of active sampling strategies when comparisons are simulated using a logistic regression model according to (2). In-sample Kendall’s Tau distance $R_{I_{D}}$ on 200 images (left) and generalization error $R_{I_{E}}-R_{I_{D}}$ for models trained on 150 images and evaluated on 150 images from a different distribution (right). All results are averaged over $100$ different random seeds.
Implementation details
GURO, CoLSTIM, and Uniform use the logistic regression implementation in Scikit-learn (Pedregosa et al., 2011) with default Ridge regularization ($C=1$). The former two update $\theta_{t}$ every iteration using the full history, $D_{t}$. The design matrix for CoLSTIM is updated as in Bengs et al. (2022): the confidence width $c_{1}$ was chosen to be $\sqrt{d\log(T)}$, and the perturbed values were generated using the standard Gumbel distribution. GURO updates $\mathbf{H}^{-1}_{t}$ every iteration using the Sherman-Morrison formula.
For BayesGURO and BALD, the posterior $p(\theta\mid D_{t})$ is estimated using the Laplace approximation as described in Bishop and Nasrabadi (2006). With this approximation, the covariance matrix is the same as the inverse of the Hessian of the log-likelihood. For both methods, the priors $\theta_{B,0}=\mathbf{0}^{d}$ and $\mathbf{H}_{B,0}^{-1}=I_{d}$ were used, and sequential updates were performed every iteration. The sample criterion for BALD under a logistic model is given in Appendix C.4.1. For BayesGURO, 50 posterior samples were used to estimate $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ for every $z_{ij}$. The hybrid algorithms follow the same structure, with a small change in sampling (see Appendix F.2), and a constraint that each per-item parameter $\zeta_{i}$ is independent of other parameters. This allows for efficient updates of $\mathbf{H}_{B,t}^{-1}$ by using sparsity in the covariance.
6.1 Ordering X-ray images under the logistic model
Our first task (X-RayAge) is to order X-ray images based on perceived age (Ieki et al., 2022) where the preference feedback follows a (well-specified) logistic model. We base this experiment on publicly available data from the Kaggle competition ”X-ray Age Prediction Challenge” Felipe Kitamura (2023), hosted by the Radiology and Diagnostic Imaging Society of São Paulo (SPR), which contains more than 10 0001000010\ 00010 000 de-identified chest X-rays (one image per subject), along with the person’s true age. Features were extracted using the 121-layer DenseNet in the TorchXrayVision package (Cohen et al., 2022) followed by PCA projection, resulting in $35$ features. A ridge regression model, $\theta_{*}$, was fit to the true age ($R^{2}\approx 0.67$). During the active learning phase, feedback is drawn from $p(C_{ij}=1)=\sigma\left(\theta_{*}^{\top}z_{i,j}\cdot\lambda\right)$ where $\lambda$ (set to 0.1) controls the noise level. We only look at the fully contextual models here since they are well-specified by design, $\mathcal{I}$ can be ordered using only contextual features.
Table 1: X-RayAge. Normalized Kendall’s Tau distance for models ordering $n=200$ X-ray images represented by $d=35$ features at different numbers of comparisons $T\in\{\log n,d,n,n\log n\}$, along with the $95\%$ CI.
$T=\log n$ $d$ $n$ $n\log n$
Uniform $0.28\pm 0.07$ $0.19\pm 0.04$ $0.12\pm 0.02$ $0.05\pm 0.01$
BALD $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.02$ $0.07\pm 0.01$
CoLSTIM $0.28\pm 0.08$ $0.19\pm 0.04$ $0.13\pm 0.02$ $0.09\pm 0.01$
NormMin $\mathbf{0.27}\pm 0.07$ $0.19\pm 0.04$ $0.15\pm 0.03$ $0.08\pm 0.01$
GURO $\mathbf{0.27}\pm 0.07$ $\mathbf{0.16}\pm 0.03$ $\mathbf{0.09}\pm 0.01$ $\mathbf{0.04}\pm 0.01$
BayesGURO $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.01$ $\mathbf{0.04}\pm 0.01$
In the first setting, we sub-sample 200 X-ray images uniformly at random from the full set. A ground-truth ordering of these elements is derived using the learned linear model. Figure 1(a) shows the ordering error over 2 000 iterations while Table 1 highlights the error for different budgets, $T$. GURO and BayesGURO perform similarly, both better than the baselines. We note that BALD starts off converging about as fast as the GURO algorithm, but plateaus, most likely as a result of actively avoiding comparisons with high aleatoric uncertainty. Finally, the poor performance of CoLSTIM highlights the discrepancy between best-arm identification and recovering a complete ordering. In Table 1, we see that NormMin performs worse than our proposed algorithms and even worse than Uniform, once the budget is large enough. This shows the importance of accounting for aleatoric uncertainty and that active sampling to optimize the second-order term in Lemma 1 is not sufficient.
In the second setting, we evaluate how well the algorithms generalize to new items. First, we sample 300 X-ray images from the full data set. Next, we split these into two sets, with one ($I_{D}$) containing the youngest $50\%$ and the other ($I_{E}$) the oldest $50\%$. The algorithms were then trained to order the list containing the younger subjects, but were simultaneously evaluated on how well they could sort the list containing the older subjects. The simulation was run for 2 00020002\ 0002 000 iterations. The continuously measured difference in ordering error evaluated on $I_{E}$ and $I_{D}$ are presented in Figure 1(b). As expected, all algorithms are worse at ordering the items in $I_{E}$, $R_{I_{E}}-R_{I_{D}}>0$. However, GURO and BayesGURO achieve the lowest average difference. Together with Figure 1(a), this means that our proposed algorithms achieve the best in-sample and out-of-sample orderings.
{subcaptionblock}
[t]0.48 Refer to caption {subcaptionblock}[t]0.48 Refer to caption
Figure 2: Kendall’s Tau distance between estimated and true ordering of 6 07260726\ 0726 072 images.
Figure 3: Ordering a list $\mathcal{I}_{D}$ of 3 00030003\ 0003 000 images. After 10 0001000010\ 00010 000 comparisons, the remaining 3 07230723\ 0723 072 images, $\mathcal{I}\setminus\mathcal{I}_{D}$, are added and $R_{\mathcal{I}}$ is plotted.
Figure 4: PhotoAge. The ordering error when comparisons are made by human annotators and no linear model is assumed. The convergence on the training images (left) as well as how well the methods generalize to new images (right). To avoid cluttering, only the Bayesian algorithms were included as they lend themselves to sequential updates.
6.2 Ordering IMDB images with human annotations
Next, we use the public IMDB Age dataset (Pavlichenko and Ustalov, 2021), to study the algorithms’ performance when preference feedback comes from human annotators (PhotoAge). The data contains 9 15091509\ 1509 150 images of actors and 250 249250249250\ 249250 249 pairwise comparisons where annotators have stated which of two people appears older. No image pairs were rated twice. After removing grayscale and low-resolution images, 6 07260726\ 0726 072 images and 110 349110349110\ 349110 349 comparisons remain. We extract features from each image using the Inception-ResNet implemented in FaceNet Schroff et al. (2015) followed by PCA, resulting in $d=75$ features per image. We generate a ground-truth ordering by sorting images by their Borda score, i.e. the fraction of comparisons where an actor was considered to be the oldest; the real ages of the actors were not used (see discussion in Appendix F.1). We restrict algorithms to only query pairs for which an annotation exists and remove the pair from the pool once queried.
In the first setting, all 6 07260726\ 0726 072 images were ordered using 50 0005000050\ 00050 000 comparisons. We compare the fully contextual (6) and hybrid (7) GURO models, and TrueSkill, a non-contextual algorithm that only uses per-item parameters, included here since the contextual models may be misspecified. The results, see Figure 4, highlight the differences between modeling strategies. The fully contextual algorithms initially converge quickly, achieving a rough ordering of the items, before plateauing and not making any real improvements. This most likely correlates with the richness of the features, meaning that while they are informative enough to roughly order the list, they are not sufficient for retrieving a more granular ordering. TrueSkill on the other hand starts by improving the ordering at a much lower pace, but keeps converging more steadily throughout. Perhaps most interesting is the result of the hybrid algorithms, which seemingly reap the benefits of both methods, converging quickly like the contextual, while not hitting a plateau. In fact, as discussed in Appendix F.3, the hybrid models perform comparably to TrueSkill even when features are completely uninformative. One can also observe that while the choice of adaptive strategy has a somewhat modest impact in the fully contextual case, it has a large effect on the hybrid models.
The second experiment is a few-shot scenario where after some time, additional images are added to the pool of items. The initial pool consists of 3 00030003\ 0003 000 images sampled from the dataset. After 10 0001000010\ 00010 000 steps, the remaining 3 07230723\ 0723 072 images were added to the pool. Figure 4 shows how the ordering error is affected by the addition of new data for all three types of models. We observe that the increase in error of the fully contextual model is very slight, most likely a result of added samples being drawn from the same distribution. In the case of TrueSkill, the error increases drastically as a result of the algorithm not having seen these items before and having no way of generalizing the results of previous comparisons to them. Lastly, the hybrid algorithms seem to be moderately affected. The error does increase as the model has not yet tuned any of the added per-item parameters, but one can also observe that the extent of this is much lower than for TrueSkill as the model can provide a rough ranking of the out-of-sample elements using the contextual features.
7 Conclusion
We have demonstrated the benefit of utilizing contextual features in active preference learning to efficiently order a list of items. Empirically, this leads to quicker convergence, compared to non-contextual methods, and allows algorithms to generalize out-of-sample. We derived an upper bound on the ordering error and used it to design two active sampling strategies that outperform baselines on realistic image ordering tasks. Both theoretical and empirical results highlight the benefit of accounting for aleatoric uncertainty when learning from human annotators. The optimality of our sampling strategies remains an open question. A future direction is to derive a lower bound on the ordering error, and prove an—ideally matching—algorithm-specific upper bound. Furthermore, for now, the benefit of utilizing item attributes depends on the quality of features that can be extracted using a pre-trained model. This could most likely be improved by performing representation learning throughout the learning process. Finally, while we perform experiments with real-world human annotations, these are limited to already-collected (offline) data. Future work should evaluate the strategies in an online setting where participants annotate comparisons recommended by the algorithms.
Appendix A Notation
Table 2: Notation
$\mathcal{I}$ Collection of items =$\mathcal{I}=\{1,...,n\}$
$n$ Number of items
$d$ Dimension of item attributes
$x_{i}\in\mathbb{R}^{d}$ Context attributes for item $i\in\mathcal{I}$
$z_{ij}\in\mathbb{R}^{d}$ $z_{ij}=x_{i}-x_{j}$ for $i,j\in\mathcal{I}$
$y_{i}$ Score for item $i\in\mathcal{I}$
$c_{t}\in\{0,1\}$ The outcome of the comparison at time $t$, $1$ if $i_{t}$ was preferred to $j_{t}$
$D_{t}$ $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$
$\theta\in\mathbb{R}^{d}$ Model parameter
$\theta_{*}\in\mathbb{R}^{d}$ Model parameter of the environment
$\theta_{t}\in\mathbb{R}^{d}$ Estimated parameter at time $t$
$\sigma(.)$ Sigmoid (logistic) function
$\dot{\sigma}(.)$ derivative of $\sigma(.)$
$\mathbf{H}_{t}(\theta)$ Hessian of the negative log-likelihood $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$
$\tilde{\mathbf{H}}_{t}(\theta)$ Hessian normalized by number of plays $\tilde{\mathbf{H}}_{t}(\theta):=\frac{1}{t}\mathbf{H}_{t}(\theta)$
$\theta_{B,t}\in\mathbb{R}^{d}$ The MAP estimate of $\theta$ at time $t$
$\mathbf{H}_{B,t}$ The Hessian in the Bayesian setting, adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$
$\|z_{ij}\|_{\mathbf{H}^{-1}_{t}(\theta)}$ $\|z_{ij}\|_{\mathbf{H}^{-1}_{t}(\theta)}=\sqrt{z_{ij}^{\top}\mathbf{H}^{-1}_{t}(\theta)z_{ij}}$
$h$ Comparison model (binary output)
$f$ Comparison logit (typically linear), e.g., $f_{\theta}(i,j)=\theta^{\top}(x_{i}-x_{j})$
Appendix B Extended related work
Preference learning aims at ordering a set of items based on (noisy) outcomes of pairwise comparisons (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005). The problem is related also to learning to rank  (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012). Adaptively collecting data to solve preference learning with as few examples as possible may be posed as an active learning (Brinker, 2004; Silva et al., 2014; Long et al., 2010; Ling et al., 2020) or bandit problem. An example of the former, Houlsby et al. (2011) proposed the algorithm and applied it to active preference learning with per-item context variables and noisy pairwise preference feedback, adaptively proposing items for comparison by maximizing the expected information gain which is a commonly used acquisition function in Bayesian optimal design (Lindley, 1956). Singh and Chakraborty (2021) study age estimation through active learning of deep neural networks, like our task in Section 6.2, but used a mix of absolute and relative feedback.
Bandit algorithms with relative feedback are often referred to as dueling bandits (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022) and have been proposed both for items with context attributes (Dudík et al., 2015) and without (Yue et al., 2012). Recently, Bengs et al. (2022) proposed CoLSTIM which learns from relative feedback in the linear stochastic transitivity model, matching the model in (2) and Di et al. (2023) gave variance-aware regret bounds for stochastic contextual dueling bandits. The bandit paradigm closest to our problem is fixed-budget pure-exploration (Fang, 2022), where the goal is to identify the highest-ranked item in a set using a fixed budget of comparisons. Related to our work is also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits. Another variant is minimization of Borda regret (Wu et al., 2023b). Jun et al. (2021) considers a logistic model with pairwise comparisons where the goal is to find the item with highest ranking and they provide guarantees in the fixed-confidence regime. However, as we argue in Appendix E, a model that identifies the highest-ranked item optimally fast can be arbitrarily bad at recovering a total ordering. Our work is also related to the generalized linear bandit (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020) and our theoretical analysis leverage techniques from this literature.
Another view of the preference learning problem is offered by non-contextual ranking methods, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018) and Probe-Rank (Lou et al., 2022). These produce an in-sample ordering of a fixed set of items but must be updated if new items are added.
We are also interested in the theoretical analysis of active preference learning. Chaudhuri et al. (2015) give an algorithm for and analysis of active learning for general maximum likelihood estimation. Ash et al. (2021) extend this idea and propose a tractable algorithm for models parameterized by deep neural network. Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information.
Preference learning is commonly used when training large language models through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Wu et al., 2023a). In this line of work, Zhu et al. (2023) is closest to our setting and they provide guarantees on the sample complexity of learning a preference model from offline data. They leverage similar tools from statistical learning and bandits as we do. In contrast to their work, we provide sampling strategies and guarantees for the online setting. Mehta et al. (2023) consider active learning for RLHF in a dueling bandit framework where the goal is to optimize a contextual version of the Borda regret.
Appendix C Algorithms
C.1 MLE estimator for logistic regression
The log-likelihood $L_{t}(\theta)$ of data $D_{t}=\{(i_{s},j_{s},c_{s})\}_{s=1}^{t}$, with $z_{s}=x_{i_{s}}-x_{j_{s}}$, under a logistic regression model with parameters $\theta$ is defined by
$L_{t}(\theta)=\sum_{s=1}^{t}\left(c_{s}\log\sigma(\theta^{\top}z_{s})+(1-c_{s})(1-\sigma(\theta^{\top}z_{s}))\right)~{}.$
The maximum likelihood estimator (MLE) at time $t$ is the parameters
$\theta_{t}=\operatorname*{arg\,max}_{\theta}L_{t}(\theta)~{}.$ (9)
The regularized estimator with ridge/$\ell_{2}$ penalty with parameter $\lambda$ is
$\theta^{R}_{t}=\operatorname*{arg\,min}_{\theta}-L_{t}(\theta)+\lambda\|\theta\|_{2}^{2}~{}.$
C.2 Bayesian estimator for logistic regression
$\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ according to the log likelihood
$\theta_{B,t}=\operatorname*{arg\,max}_{\theta}\ln p(\theta\mid D_{t}),$ (10)
where
$\begin{split}\ln p(\theta\mid D_{t})=&-\frac{1}{2}(\theta-\theta_{B,0})^{\top}\mathbf{H}^{-1}_{B,0}(\theta-\theta_{B,0})\\ &+\sum_{t}c_{t}\ln(\sigma(z_{i_{t},j_{t}}^{\top}\theta))+(1-c_{t})\ln(1-\sigma(z_{i_{t},j_{t}}^{\top}\theta))+const.\end{split}$
The hessian at time $t$ is defined as
$\mathbf{H}_{B,t}=\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}=\mathbf{H}_{B,0}+\mathbf{H}_{t}.$
Note that if priors $\theta_{B,0}=\mathbf{0}$ and $\mathbf{H}^{-1}_{B,0}=I_{d}$ are used, the log likelihood boils down to:
$\begin{split}\ln p(\theta\mid D_{t})=&-\frac{1}{2}||\theta||_{2}^{2}+\sum_{t}c_{t}\ln(\sigma(z_{i_{t},j_{t}}^{\top}\theta))+(1-c_{t})\ln(1-\sigma(z_{i_{t},j_{t}}^{\top}\theta))+const\end{split}$
which implies that the MAP estimate will be the same as the MLE estimate in the frequentist setting with a ridge regularisation. Similarly the Hessian becomes:
$\mathbf{H}_{B,t}=\mathbf{H}_{t}+I_{d}$
C.3 Uniform sampling
The uniform sampling algorithm is given in Algorithm 2. The corresponding Bayesian version replaces line 5 with the MAP estimate.
Algorithm 2 Uniform sampling algorithm
1:Training items $\mathcal{I}_{D}$, attributes ${\bf X}=\{x_{i}\}_{i\in\mathcal{I}_{d}}$
2:for $t=1,...,T$ do
3:    Sample $(i_{t},j_{t})$ uniformly
4:    Observe $c_{t}$ from noisy comparison (annotator)
5:    $D_{t}=D_{t-1}\cup\{i_{t},j_{t},c_{t})\}$
6:    Let $\theta_{t}=\mathrm{MLE}(D_{t})$
7:end for
8:Return $h_{T}$
C.4 BALD
Algorithm 3 BALD bandit
1:Training items $\mathcal{I}_{D}$, attributes ${\bf X}=\{x_{i}\}_{i\in\mathcal{I}_{d}}$
2:Initialize $\theta_{B,0}=\mathbf{0}$, $\mathbf{H}_{B,0}=\lambda^{-1}I$
3:for $t=1,...,T$ do
4:    Draw $(i_{t},j_{t})$ = $\operatorname*{arg\,max}_{i,j}H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$
5:    Observe $c_{t}$ from noisy comparison (annotator)
6:    $D_{t}=D_{t-1}\cup\{i_{t},j_{t},c_{t})\}$
7:    Let $\theta_{t}=\mathrm{MAP}(D_{t})$
8:    Update ${\bf H}_{B,t}\leftarrow\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta_{t})z_{i,j}z_{i,j}^{\top}$
9:end for
10:Return $h_{T}$
Where the posterior is calculated as in Appendix C.2 and $H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ is approximated as in Appendix C.4.1.
C.4.1 Deriving the BALD sampling criterion
The BALD criteria formalized using our notation becomes
$\operatorname*{arg\,max}_{i,j}H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]],$
where $H$ represents Shannon’s entropy
$h(p)=-p\log_{2}(p)-(1-p)\log_{2}(1-p).$
The first term of the equation becomes
[…]
Here $\Pr(y\mid z_{i,j},D_{t})$ is the predictive distribution for our Bayesian logistic regression model. As covered in Bishop and Nasrabadi (2006), this expectation cannot be evaluated analytically but can be approximated using the probit function $\Phi$;
$\Pr(y\mid z_{ij},D_{t})\approx\Phi\left(\frac{\mathbf{\theta}^{\top}_{t}z_{i,j}}{\sqrt{\lambda^{-2}+||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\approx\sigma\left(\frac{\mathbf{\theta}^{\top}_{t}z_{i,j}}{\sqrt{1+\frac{\pi||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}(\theta_{*})}}{8}}}\right).$
Next, the term $\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ must be calculated. The true definition is
[…]
To make this a one variable integral, let $X=\theta^{\top}z_{i,j}$ define a new random variable. Since $\theta\sim\mathcal{N}(\theta_{t},\mathbf{H}_{t}^{-1})$, and $z_{i,j}$ is just a constant vector, we know that $X$ will follow a univariate normal distribution $X\sim\mathcal{N}(\theta_{t}^{\top}z_{i,j},||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}})$. This allows us to rewrite the integral as
[…]
However, this integral has no closed form solution. Instead we perform the same strategy as in Houlsby et al. (2011) and do a Taylor expansion of $\ln{h(\sigma(\mathbf{\theta}^{\top}\mathbf{z}))}$. The third-order Taylor expansion gives us
$h(\sigma(x))\approx\exp{\left(-\frac{x^{2}}{8\ln 2}\right)}.$
Inserting this, the term can be approximated as
[…]
where $C=\sqrt{4\ln{2}}$. Finally, we arrive at an estimation of the objective function which we wish to maximize:
$H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]\approx h\left(\sigma\left(\frac{\theta_{t}^{\top}z_{i,j}}{\sqrt{1+\frac{\pi}{8}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\right)-\frac{C}{\sqrt{||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2}}}\exp{\left(-\frac{(\theta_{t}^{\top}z_{i,j})^{2}}{(||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2})}\right)}$
Appendix D Proofs of Lemma 1 and Theorem 1
D.1 Proof of Lemma 1
Proof.
We now proceed to bound
$P\left(|\sigma(z_{ij}^{\top}\theta_{t})-\sigma(z_{ij}^{\top}\theta_{*})|>\Delta\right).$
From the self-concordant property of logistic regression we have (Faury et al., 2020)
$|\sigma(z_{ij}^{\top}\theta_{t})-\sigma(z_{ij}^{\top}\theta_{*})|\leq\dot{\sigma}(z_{ij}\top\theta_{t})|z_{ij}^{\top}(\theta_{t}-\theta_{*})|+\frac{1}{4}|z_{ij}^{\top}(\theta_{t}-\theta_{*})|^{2}.$
We will prove a high probability bound on the event
$\dot{\sigma}(z_{ij}\top\theta_{t})|z_{ij}^{\top}(\theta_{t}-\theta_{*})|+\frac{1}{4}|z_{ij}^{\top}(\theta_{t}-\theta_{*})|^{2}\leq\Delta.$ (11)
Directly trying to bound the LHS in Equation 11 will result in a rather messy expression. Instead, we define the events
$\mathcal{E}_{1}$ :=$:=\left\{\dot{\sigma}(z_{ij}\top\theta_{t})|z_{ij}^{\top}(\theta_{t}-\theta_{*})|\leq\frac{\Delta}{2}\right\}$
$\mathcal{E}_{2}$ :=$:=\left\{\frac{1}{4}|z_{ij}^{\top}(\theta_{t}-\theta_{*})|^{2}\leq\frac{\Delta}{2}\right\}.$
Clearly $\mathcal{E}_{1}\bigcup\mathcal{E}_{2}$ implies the expression in Equation 11. Assume we have bounds on the complement of these events, $P\left(\mathcal{E}_{1}^{c}\right)\leq\alpha$ and $P\left(\mathcal{E}_{2}^{c}\right)\leq\beta$. Then
$P\left(|\sigma(z_{ij}^{\top}\theta_{t})-\sigma(z_{ij}^{\top}\theta_{*})|>\Delta\right)$ $\leq\alpha+\beta+\alpha\beta$
$\leq 2\alpha+2\beta.$
We now proceed to bound the probability of these complements separately.
Step 1. Relating $\theta_{t}$ to $\theta_{*}$: The first challenge in our analysis to is relate $\theta_{*}$ and $\theta_{t}$. In contrast to linear regression, where we have a closed-form expression for $\theta_{t}$, there is no analytical solution for $\theta_{t}$ given a set of observation. However, we know that $\theta_{t}$ is the MLE, corresponding to
$\theta_{t}=\operatorname*{arg\,max}_{\theta}L_{t}(\theta)$
where
$L_{t}(\theta)=\sum_{s=1}^{t}c_{s}\log\sigma\left(z_{s}^{\top}\theta\right)+(1-c_{s})\log\left(1-\sigma\left(z_{s}^{\top}\theta\right)\right).$
We have
$\nabla_{\theta}L_{t}(\theta)=\sum_{s=1}^{t}c_{s}z_{s}-\underbrace{\sum_{s=1}^{t}\sigma\left(z_{s}^{\top}\theta\right)z_{s}}_{g_{t}(\theta)}$
and hence $g_{t}(\theta_{t})=\sum_{s=1}^{t}c_{s}z_{s}$.
A standard trick in logistic bandits (Filippi et al., 2010; Faury et al., 2020; Jun et al., 2021) is to relate $\theta_{*}-\theta_{t}$ to $g_{t}(\theta_{*})-g_{t}(\theta_{t})$. Especially, the following equality is due to the mean-value theorem (see Filippi et al. (2010))
$g_{t}(\theta_{*})-g_{t}(\theta_{t})=\mathbf{H}_{t}(\theta^{\prime})\left(\theta_{*}-\theta_{t}\right)$ (12)
where $\theta^{\prime}$ is some convex combination of $\theta_{*},\theta_{t}$. Note that $\mathbf{H}_{t}(\theta^{\prime})$ has full rank.
Using Equation 12 yields
$\left|z_{ij}^{\top}\left(\theta_{*}-\theta_{t}\right)\right|$ $=\left|z_{ij}^{\top}\mathbf{H}^{-1}_{t}(\theta^{\prime})\left(g_{t}(\theta_{*})-g_{t}(\theta_{t})\right)\right|$
Furthermore, since $g_{t}(\theta_{t})=\sum_{s=1}^{t}c_{s}z_{s}$, due to $\nabla_{\theta}L_{t}(\theta_{t})=0$, we have
$g_{t}(\theta_{t})-g_{t}(\theta_{*})=\sum_{s=1}^{t}\underbrace{\left(c_{s}-\sigma\left(z_{s}^{\top}\theta_{*}\right)\right)}_{\epsilon_{s}}z_{s}$
where $\epsilon_{s}$ is a sub-Gaussian random variable with mean 00 and variance $\nu_{s}^{2}:=\dot{\sigma}\left(z_{s}^{\top}\theta_{*}\right)$. We define
$S_{t}:=\sum_{s=1}^{t}\epsilon_{s}z_{s}.$
We now have
$\left|z_{ij}^{\top}\left(\theta_{*}-\theta_{t}\right)\right|$ $=\left|z_{ij}^{\top}\mathbf{H}^{-1}_{t}(\theta^{\prime})S_{t}\right|$
and Lemma 10 in Faury et al. (2020) states that $\mathbf{H}_{t}^{-1}(\theta^{\prime})\preccurlyeq(1+2S)\mathbf{H}_{t}^{-1}(\theta_{*})$ where $||\theta_{*}||_{2}\leq S$. Hence,
$\left|z_{ij}^{\top}\left(\theta_{*}-\theta_{t}\right)\right|\leq\left(1+2S\right)\left|z_{ij}^{\top}\mathbf{H}^{-1}_{t}(\theta_{*})S_{t}\right|$
Step 2. Tail bound for vector-valued martingales:
We will now prove an upper bound on the probability that $\left|z_{ij}^{\top}\mathbf{H}^{-1}_{t}(\theta_{*})S_{t}\right|$ deviates much from a certain threshold. This step is based on the proof of Lemma 1 in Filippi et al. (2010) which itself is based on a derivation of a concentration inequality in Rusmevichientong and Tsitsiklis (2010). The difference compared to Filippi et al. (2010) is that we work with the Hessian $\mathbf{H}_{t}(\theta_{*})$ instead of the design matrix for linear regression $V_{t}=\sum_{s}x_{s}x_{s}^{\top}$. This require us to construct a slightly different martingale.
Let $A$ and $B$ are two random variables such that
$\mathbb{E}\left[\exp\left\{\gamma A-\frac{\gamma^{2}}{2}B^{2}\right\}\right]\leq 1,\forall\gamma\in\mathbb{R}$ (13)
then due to Corollary 2.2 in de la Peña et al. (2004) it holds that $\forall a\geq\sqrt{2}$ and $b>0$
$P\left(|A|\geq a\sqrt{(B^{2}+b)\left(1+\frac{1}{2}\log\left(\frac{B^{2}}{b}+1\right)\right)}\right)\leq\exp\left\{\frac{-a^{2}}{2}\right\}.$ (14)
Let $\eta\in\mathbb{R}^{d}$ and consider the process
$M^{\gamma}_{t}(\theta_{*},\eta):=\exp\left\{\gamma\eta^{\top}S_{t}-\gamma^{2}||\eta||^{2}_{\mathbf{H}_{t}(\theta_{*})}\right\}.$ (15)
We will now proceed to prove that $M^{\gamma}_{t}(\theta,\eta)$ is a non-negative super martingale satisfying Equation 13. Note that
$\gamma\eta^{\top}S_{t}-\gamma^{2}||\eta||^{2}_{\mathbf{H}_{t}(\theta_{*})}=\sum_{s=1}^{t}\underbrace{\left(\gamma\eta^{\top}z_{s}\epsilon_{s}-\dot{\sigma}(\theta^{\top}z_{s})\gamma^{2}\left(\eta^{\top}z_{s}\right)^{2}\right)}_{F_{s}}=\sum_{s=1}^{t}F_{s}.$
Further we use the fact that $\epsilon_{s}$ is sub-Gaussian with parameter $\nu_{s}$, .i.e,
$\mathbb{E}\left[\exp\{\lambda\epsilon_{s}\}\right]\leq\exp\left\{\nu_{s}^{2}\lambda^{2}\right\},\forall\lambda>0.$
Let $D_{s-1}$ denote the observations up until time $s$, then
$\mathbb{E}\left[\exp\{F_{s}\}\mid D_{s-1}\right]$ $=\mathbb{E}\left[\exp\left\{\underbrace{\gamma\eta^{\top}z_{s}}_{\lambda}\epsilon_{s}\right\}\right]\exp\left\{-\underbrace{\dot{\sigma}(\theta_{t}^{\top}z_{s})}_{\nu_{s}^{2}}\gamma^{2}\left(\eta^{\top}z_{s}\right)^{2}\right\}$
$\leq\exp\left\{\nu_{s}^{2}\lambda^{2}\right\}\exp\left\{-\nu_{s}^{2}\lambda^{2}\right\}=1.$
This also implies
$\mathbb{E}\left[M^{\gamma}_{t}(\theta_{*},\eta)\mid D_{t-1}\right]\leq M^{\gamma}_{t-1}(\theta_{*},\eta)$
and $M^{\gamma}_{t}(\theta_{*},\eta)$ is a super-martingale satisfying
$\mathbb{E}\left[\exp\left\{\gamma\eta^{\top}S_{t}-\gamma^{2}||\eta||^{2}_{\mathbf{H}_{t}(\theta_{*})}\right\}\right]\leq 1,\forall\gamma\geq 0$
and we can apply the results of de la Peña et al. (2004).
We now follow the last step of the proof of Lemma 1 in Filippi et al. (2010). We let $a=\sqrt{2\log\frac{1}{\delta}}$ for some $\delta\in(0,1/e)$ and let $b=\lambda_{0}\|\eta\|^{2}_{2}$. We have with probability at least $1-\delta$
$|\eta^{\top}S_{t}|\leq\sqrt{2\log\frac{1}{\delta}}\sqrt{\|\eta\|^{2}_{\mathbf{H}_{t}(\theta_{*})+\lambda_{0}\|\eta\|^{2}_{2}}\left(1+\frac{1}{2}\log\left(1+\frac{\|\eta\|^{2}_{\mathbf{H}_{t}(\theta_{*})}}{\lambda_{0}\|\eta\|^{2}_{2}}\right)\right)}.$
Rearanging and using the fact that $\lambda_{0}||\eta||^{2}_{2}\leq\|\eta\|^{2}_{\mathbf{H}_{t}(\theta_{*})}\leq t\|\eta\|_{2}$ yields
$|\eta^{\top}S_{t}|\leq\rho(\lambda_{0})||\eta||_{\mathbf{H}_{t}(\theta_{*})}\sqrt{2\log\frac{t}{\delta}}.$ (16)
where $\rho$ is defined as
$\rho(\lambda_{0})=\sqrt{3+2\log\left(1+\frac{4Q^{2}}{\lambda_{0}}\right)}.$
We take $M_{t}$ to be a matrix such that $M_{t}^{2}=\mathbf{H}_{t}(\theta_{*})$ and note that for any $\tau>0$
$P\left(||S_{t}||_{\mathbf{H}^{-1}_{t}(\theta_{*})}^{2}\geq d\tau^{2}\right)\leq\sum_{i=1}^{d}P\left(\left|S_{t}^{\top}M_{t}^{-1}e_{i}\right|\geq\tau\right)$
where $e_{i}$ is the i:th unit vector. Equation 16 with $\eta=M_{t}^{-1}e_{i}$ together with $||M_{t}^{-1}e_{i}||_{\mathbf{H}_{t}(\theta_{*})}=1$ yield that the following holds with with probability at least $1-\delta$
$||S_{t}||_{\mathbf{H}_{t}^{-1}(\theta_{*})}\leq\rho(\lambda_{0})\sqrt{2d\log t}\sqrt{\log\frac{d}{\delta}}.$ (17)
Step 3. (Unverifiable) High-probability bounds on $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$.
We now have enough machinery to state high-probability bounds for our two events. These bounds will be unverifiable in the sense that the depend on the true parameter $\theta_{*}$ which is not known to us during runtime. We derive verifiable bounds in the next step of the proof.
Recall that $\mathbf{H}_{t}^{-1}(\theta_{*})$ is symmetric. We apply Equation 16 with $\eta=\mathbf{H}_{t}^{-1}(\theta_{*})z_{ij}$ and $\alpha>0$ in place of $\delta$. First, we note that $\|\mathbf{H}_{t}^{-1}(\theta_{*})z_{ij}\|_{\mathbf{H}_{t}(\theta_{*})}=\|z_{ij}\|_{\mathbf{H}^{-1}_{t}(\theta_{*})}$ which implies with probability at least $1-\alpha$
$\left|z_{ij}^{\top}\mathbf{H}_{t}^{-1}(\theta_{*})S_{t}\right|=\left|S_{t}^{\top}\mathbf{H}_{t}^{-1}(\theta_{*})z_{ij}\right|\leq\rho(\lambda_{0})\|z_{ij}\|_{\mathbf{H}^{-1}_{t}(\theta_{*})}\sqrt{2\log\frac{t}{\alpha}}.$ (18)
We solve for smallest possible $\alpha\in(0,1/e)$ such that
$(1+2S)\rho(\lambda_{0})\dot{\sigma}(z_{ij}^{\top}\theta_{*})||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{*})}\sqrt{2\log\frac{t}{\alpha}}\leq\frac{\Delta}{2}$
Rearanging yields
$\alpha\leq\exp\left\{\frac{-\Delta^{2}}{8\rho^{2}(\lambda_{0})(1+2S)^{2}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{*})||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{*}))}\right)^{2}}+\log T\right\}.$ (19)
For $\mathcal{E}_{2}$ and the bound on its probability, $\beta>0$ we have
$\frac{1}{4}|z_{ij}^{\top}(\theta_{t}-\theta_{*})|^{2}\leq\frac{1}{2}(1+2S)^{2}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}(\theta_{*})}\rho^{2}(\lambda_{0})\log\frac{t}{\beta}\leq\frac{\Delta}{2}$
and
$\beta\leq\exp\left\{\frac{-\Delta}{\rho^{2}(\lambda_{0})(1+2S)^{2}\left(||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{*}))}\right)^{2}}+\log T\right\}.$ (20)
Note that both Equation 19 and Equation 20 are under the assumption that the RHS satisfy $<1/e$ since this is required in order to apply the results of de la Peña et al. (2004). As we discuss in the main text, these quantities are approaching zero as $O(Te^{-T})$, ignoring various constants, for reasonable sampling strategies and will satisfy this condition eventually.
Step 4. (Verifiable) High-probability bounds on $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$.
The bounds in the previous step depend on the true parameter $\theta_{*}$ which we do not have access to in practise. We again use Lemma 10 of Faury et al. (2020) together with Cauchy-Schwartz
$|z_{ij}(\theta_{*}-\theta_{t})|$ $=\left|z_{ij}^{\top}\mathbf{H}^{-1/2}_{t}(\theta^{\prime})\mathbf{H}^{1/2}_{t}(\theta^{\prime})S_{t}\right|$
$\leq(1+2S)||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{t})}||S_{t}||_{\mathbf{H}^{-1}_{t}(\theta_{*})}.$
Using Equation 17 we have with probability at last $1-\alpha$
$(1+2S)\sigma(z_{ij}^{\top}\theta_{*})||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{t})}||S_{t}||_{\mathbf{H}^{-1}_{t}(\theta_{*})}\leq(1+2S)\dot{\sigma}(z_{ij}^{\top}\theta_{*})||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{t})}\rho(\lambda_{0})\sqrt{2d\log t}\sqrt{\log\frac{d}{\alpha}}.$ (21)
We solve for smallest $\alpha\in(1/e)$ such that Equation 21 is smaller than $\Delta_{ij}/2$. This yields
$\alpha\leq\exp\left\{\frac{-\Delta^{2}}{8d\rho^{2}(\lambda_{0})(1+2S)^{2}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{T})||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{T}))}\right)^{2}}+\log dT\right\}.$
Same steps for $\beta$ yields
$\beta\leq\exp\left\{\frac{-\Delta}{d\rho^{2}(\lambda_{0})(1+2S)^{2}\left(||z_{ij}||_{\mathbf{H}^{-1}_{t}(\theta_{T}))}\right)^{2}}+\log dT\right\}.$
For brevity, define $C_{1}=\rho^{2}(\lambda_{0})(1+2S)^{2}$.
Using the definition of $\tilde{\mathbf{H}}_{t}$ yields the statement of Lemma 1. ∎
D.2 Proof of Theorem 1
Proof.
We let $i\succ j$ denote that $i$ is preferred to $j$. W.l.o.g assume $1\succ 2\succ...\succ n$. The key observation is that the assumption $\sigma(z_{ij}^{\top}\theta_{*})-0.5>\Delta_{ij}>\Delta_{*},\forall i,j$ implies that for any $i$ and $j$ such that $i<j$ it holds that
$\Delta_{i,j}>(j-i)\Delta_{*}.$
If we get the wrong relation between $i,j$ then $\sigma(z_{ij}^{\top}\theta_{*})-\sigma(z_{ij}^{\top}\theta_{T})>(j-i)\Delta_{*}$. Lemma 1 implies
$P(\sigma(z_{ij}^{\top}\theta_{*})-$ $\sigma(z_{ij}^{\top}\theta_{T})>(j-i)\Delta)\leq dT(\underbrace{\exp\left\{\frac{-(j-i)\Delta^{2}}{8d\rho^{2}(\lambda_{0})(1+2S)^{2}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{T})||z_{ij}||_{H^{-1}_{t}(\theta_{T}))}\right)^{2}}\right\}}_{\alpha_{ij}^{j-i}}$
$+\underbrace{\exp\left\{\frac{-(j-i)\Delta}{d\rho(\lambda_{0})(1+2S)^{2}\left(||z_{ij}||_{H^{-1}_{t}(\theta_{T}))}\right)^{2}}\right\}}_{\beta_{ij}^{j-i}}).$
Let $R(\theta_{T})$ be the ordering error of the $n$ items. Then, under a uniform distribution over items we have
$\mathbb{E}[R(\theta_{T})]$ $\leq\frac{4dT}{n(n-1)}\left(\underbrace{\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\alpha_{ij}^{j-i}}_{A}+\underbrace{\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\beta_{ij}^{j-i}}_{B}\right)$ (22)
$A$ and $B$ will be upper bounded using the same argument. We now upper bound sum $A$
Let $\alpha_{*}:=\exp\left\{\frac{-\Delta_{*}^{2}}{8dC_{1}\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}(\theta_{*})}}\right\}$ then
$A$ $\leq\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\alpha_{*}^{j-i}\leq(n-1)\left(\sum_{j=0}^{n}\alpha_{*}^{j}-1\right)$
$\leq(n-1)\left(\frac{1}{1-\alpha_{*}}-1\right).$
This follows from the definition of $\delta_{1,*}$ and properties of the geometric sum. It is easy to see that $\frac{1}{1-e^{-x}}-1=\frac{1}{e^{x}-1}$. Hence,
$\frac{4dT}{n(n-1)}A\leq\frac{4dT}{n}\left(\alpha_{*}^{-1}-1\right)^{-1}.$
For $B$ we perform the same steps with $\beta_{*}:=\exp\left\{\frac{-\Delta_{*}}{dC_{1}\max_{i,j}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}(\theta_{*})}}\right\}$ to get
$\frac{4dT}{n(n-1)}A\leq\frac{4dT}{n}\left(\beta_{*}^{-1}-1\right)^{-1}.$
Combing yields and
$\mathbb{E}\left[R(\theta_{T})\right]$ $\leq\frac{4dT}{n}\left(\left(\alpha_{*}^{-1}-1\right)^{-1}+\left(\beta_{*}^{-1}-1\right)^{-1}\right)$
By Markov’s inequality we have
$P(R(\theta_{T})\geq\epsilon)\leq\frac{4dT}{\epsilon n}\left(\left(\alpha_{*}^{-1}-1\right)^{-1}+\left(\beta_{*}^{-1}-1\right)^{-1}\right).$ (23)
Appendix E Comparison with regret minimization
Bengs et al. (2022) considered a problem formulation where the goal is to learn a parameter $\theta$ which determines the utility $Y_{i,t}$ for a set of arms $i=1,...,n$ as a function of observed context vectors $x_{i,t}$ in a sequence of rounds $t=1,...,T$,
$Y_{i,t}=\theta^{\top}X_{i,t}~{}.$
The probability that item $i$ is preferred over $j$ (denoted $i\succ j$) in round $t$ is decided through a comparison function $F$,
$\Pr(i\succ j\mid X_{i,t},X_{j,t})=F(Y_{i,t}-Y_{j,t})~{}.$
The goal in their setting is to, in each round, select two items $(i_{t},j_{t})$ so that their maximum (or average) utility is as close as possible to the utility of the best item. The expected regret in their average-utility setting is
$\Re_{BSH}=\mathbb{E}[\sum_{t=1}^{T}2Y_{i_{t}^{*},t}-Y_{i_{t},t}-Y_{j_{t},t}]~{}.$
Proposition 1 (Informal).
An algorithm which achieves minimal regret in the setting of Bengs et al. (2022) can perform arbitrarily poorly in our setting.
Proof.
The optimal choice of arm pair in the BSH setting is the optimal and next-optimal arm $(i_{t}^{*},i^{\prime}_{t})$ such that $i_{t}^{*}\succ i^{\prime}_{t}\succ j$ for any other arms $j$. Assume that the ordering of all other arms $j$ is determined by a feature $X_{j,t}(k)$ but that $X_{i_{t}^{*},t}(k)=X_{i^{\prime}_{t},t}(k)$. Then, no knowledge will be gained about arms other than the top 2 choices under the BSH regret. As the number of arms grows larger, the error in our setting grows as well. ∎
Saha (2021) study the same average-utility regret setting and give a lower bound under Gumbel noise. Saha and Krishnamurthy (2022) investigated where there is a computationally efficient algorithm that achieves the derived optimality guarantee.
Appendix F Experiment details
F.1 Ground truth on PhotoAge
While the IMDB dataset used in the PhotoAge task contains an associated age with every image, these were not used. There were two reasons for this. First, some of these ages were incorrectly assigned (often in photos containing multiple people that had been cropped to one individual), causing some noise. Second, and most importantly, we are interested in the subjective assessment of how old someone appears to be, which does not have to be consistent with the true age. The consequence of this is that if you keep collecting comparisons, there is no guarantee that the ordering will ever converge to reflect the true age. The rationale behind using the Borda Score instead is that the simple algorithm does not make any assumptions regarding transitivity Nihar B. Shah (2017), thus not favoring any of the sampling strategies.
F.2 Implementation Details PhotoAge
To increase computational efficiency, the hybrid algorithms did not evaluate all 100 000similar-toabsent100000$\sim$ 100\ 000∼ 100 000 comparisons at every time step. Instead, a subset of 5 00050005\ 0005 000 comparisons was first sampled, and the highest scoring pair in this set was chosen. This resulted in a large speed up, and no noticeable change in performance during evaluation.
F.3 Additional figures
NormMin Comparison
To highlight the importance of the first-order term in Lemma 1, we evaluated NormMin on the same X-ray ordering task as in Figure 1(a). The results, shown in Figure 5, indicate that not only does the algorithm perform worse than GURO, but is seemingly also outperformed by a uniform sampling strategy.
Figure 5: X-RayAge. NormMin included in the experiment shown in Figure. 1(a)
Synthetic Example and Illustration of Upper Bound
In this setting, $100$ synthetic data points were generated. Each data point consisted of $10$ features, where the feature values were sampled according to a standard normal distribution. The true model, $\theta_{*}$, was generated by sampling each value uniformly between $-3$ and $3$. The pairwise comparison feedback was simulated the same way as in Section 6.1, with $\lambda=0.5$. The upper bound of the probability that $R(\theta_{t})\geq 0.2$ was calculated every iteration according to Theorem 1. Each algorithm was run for 2000 comparisons, updating every 10th, the results of which can be seen in Figure 6. We observe in Figure 6(b) that our greedy algorithms are seemingly the fastest at minimizing the upper bound. The order of performance follows the same trend as in the experiments of Section 6.
(a) The risk $R(\theta_{t})$, defined as the normalized Kendall’s tau distance between estimated and true orderings.
(b) The probability that the frequency of pairwise inversions is $\geq$ $20\%$ after every comparison, according to (1).
Figure 6: The loss (left) along with the upper bound (right) when ordering a list of size 100 in a synthetic environment. The results have been averaged over 50 seeds.
Randomly initialized representation
As discussed in Section 6.2, the performance of our contextual approach will depend on the quality of the representations. To underscore the practical usefulness of our algorithms, we have performed the same experiment as in Figure 4, but this time the model used to extract image features was untrained (i.e. the weights were random). As to be expected, the results, shown in Figure 7, demonstrate that the fully contextual algorithms have no real way of ordering the items according to these uninformative features. However, the hybrid BayesGURO algorithm performs similarly to TrueSkill, despite model misspecification. This is promising, since you may not know in advance how informative the extracted features will be for the target ordering task.
Figure 7: PhotoAge. The same experiment as presented in Figure. 4, but the model used for feature extraction is untrained.

### v2

@@
 Abstract
-Learning an ordering of items based on noisy pairwise comparisons is useful when item-specific labels are difficult to assign, for example, when annotators have to make subjective assessments.
-Algorithms have been proposed for actively sampling comparisons of items to minimize the number of annotations necessary for learning an accurate ordering.
-However, many ignore shared structure between items, treating them as unrelated, limiting sample efficiency and precluding generalization to new items.
-In this work, we study active learning with pairwise preference feedback for ordering items with contextual attributes, both in- and out-of-sample.
-We give an upper bound on the expected ordering error incurred by active learning strategies under a logistic preference model, in terms of the aleatoric and epistemic uncertainty in comparisons, and propose two algorithms designed to greedily minimize this bound.
-We evaluate these algorithms in two realistic image ordering tasks, including one with comparisons made by human annotators, and demonstrate superior sample efficiency compared to non-contextual ranking approaches and active preference learning baselines.
+Learning an ordering of items based on pairwise comparisons is useful when items are difficult to rate consistently on an absolute scale, for example, when annotators have to make subjective assessments.
+When exhaustive comparison is infeasible, actively sampling item pairs can reduce the number of annotations necessary for learning an accurate ordering.
+However, many algorithms ignore shared structure between items, limiting their sample efficiency and precluding generalization to new items.
+It is also common to disregard how noise in comparisons varies between item pairs, despite it being informative of item similarity.
+In this work, we study active preference learning for ordering items with contextual attributes, both in- and out-of-sample.
+We give an upper bound on the expected ordering error of a logistic preference model as a function of which items have been compared.
+Next, we propose an active learning strategy that samples items to minimize this bound by accounting for aleatoric and epistemic uncertainty in comparisons.
+We evaluate the resulting algorithm, and a variant aimed at reducing model misspecification, in multiple realistic ordering tasks with comparisons made by human annotators.
+Our results demonstrate superior sample efficiency and generalization compared to non-contextual ranking approaches and active preference learning baselines.
 1 Introduction
 The success of supervised learning is built on annotating items at great volumes with small error.
-For subjective assessments, however, assigning a value from an arbitrary rating scale can be difficult, causing many to favor preference feedback from pairwise comparisons  (Phelps et al., 2015; Yannakakis and Martínez, 2015; Christiano et al., 2017; Zhu et al., 2023).
+For subjective assessments, however, assigning a value from an arbitrary rating scale can be difficult and prone to inconsistencies, causing many to favor preference feedback from pairwise comparisons  (Yannakakis and Martínez, 2015; Christiano et al., 2017; Ouyang et al., 2022; Zhu et al., 2023).
 Preference feedback is sufficient to learn an ordering of items (Fürnkranz and Hüllermeier, 2003), but for $n$ items, there are $\mathcal{O}(n^{2})$ possible pairs of items to compare.
-A common solution is to use crowd-sourcing  (Chen et al., 2013; Yang et al., 2021; Larkin et al., 2022), but what if the task requires certain expertise and annotations are expensive to collect?
-This is the case in the field of medical imaging, where annotations require trained radiologists (Phelps et al., 2015; Jang et al., 2022; Lidén et al., 2024, Appendix A).
-To be cost-efficient, we must select comparisons carefully, to minimize the annotations needed to learn a good ordering.
-There is a large literature on ordering a set of items $\mathcal{I}$ using only noisy pairwise comparisons of actively chosen item pairs (Herbrich et al., 2006; Maystre and Grossglauser, 2017; Heckel et al., 2018; Lou et al., 2022).
-While robust against noise and more efficient than exhaustive comparison, these algorithms ignore any context attributes that the items may have.
-Say that a radiologist wants to quantify the expression of a disease in a collection of X-ray images $\mathcal{I}$.
-Ranking these images according to their severity $y_{i}$ using purely preference-based algorithms ignores the content $x_{i}$ of the X-rays, sampling images $i,j\in\mathcal{I}$ for comparison based only on the current estimated severity and previous feedback.
-However, the content of the images reveal similarities between X-rays which can inform an ordering strategy.
-In applications, the set $\mathcal{I}$ we want to order is often larger than the set of items we observe during training—we may want to rank new images in relation to previous ones.
-However, methods based only on preference feedback cannot generalize what they learned to new items.
-As an alternative, active learning algorithms for classification can be used for ordering by fitting a map from pairs of contexts $x_{i},x_{j}$ to the comparison $i\succ_{?}j$, that can be applied to new items (Houlsby et al., 2011; Kirsch et al., 2019; Bengs et al., 2022).
-However, to the best of our knowledge, there is no analysis that justifies using these algorithms to learn a complete ordering or gives guarantees for their performance.
+A common solution is to use crowd-sourcing  (Chen et al., 2013; Yang et al., 2021; Larkin et al., 2022), but many tasks require domain expertise, making annotations expensive to collect.
+This is the case in the field of medical imaging, where annotations require trained radiologists (Phelps et al., 2015; Jang et al., 2022; Lidén et al., 2024; Tärnåsen and Bergström, 2023).
+So, how can we learn the best ordering possible from a limited number of comparisons?
+Classically, this problem is solved by active learning, sampling comparisons based on preference feedback and estimated item scores (Herbrich et al., 2006; Maystre and Grossglauser, 2017; Heckel et al., 2018).
+However, consider a radiologist who wants to quantify the expression of a disease in a collection of X-ray images.
+Purely preference-based algorithms utilize only the outcomes of comparisons but ignore the contents of the X-rays, which can reveal similarities between items and inform an ordering strategy.
+Moreover, the set we want to order is often larger than the set of items observed during training—we may want to rank new X-rays in relation to previous ones.
+This cannot be solved by learning per-item scores alone.
+As an alternative, active learning for classification can be used to fit a map from pairs of item contexts $x_{i},x_{j}$ (e.g., the contents of images) to the comparison $i\succ_{?}j$, that can be applied to old and new items alike (Houlsby et al., 2011; Qian et al., 2015).
+However, as we show in Section 4, learning this map to recover a complete ordering is distinct from the tasks preference learning is commonly used for, and existing algorithms lack theoretical justification for this application.
+Moreover, formal results for related problems, such as contextual bandits or reinforcement learning (Das et al., 2024; Filippi et al., 2010; Zhu et al., 2023; Bengs et al., 2022), do not translate directly to effective active sampling criteria for ordering.
+There is a small body of work on learning a contextual model to recover the complete ordering (Jamieson and Nowak, 2011; Ailon, 2011) but these either assume noiseless preference feedback or that the noise is unrelated to the similarity of items, which is unrealistic for subjective assessments.
 Contributions.
-In this work, we study active preference learning algorithms for ordering sets of items based on noisy pairwise comparisons.
-We propose the use of a contextual logistic preference model to support efficient ordering in-sample as well as generalization to new items.
-Our analysis yields the first bound on the expected ordering error achievable under this model, provided a fixed budget of comparisons (Section 4).
-This result justifies an active sampling principle based on the epistemic and aleatoric uncertainty in the outcome of a comparison between two items.
-We use this in one greedy, deterministic, and one Bayesian, stochastic algorithm (Section 5) applied to both a fully contextual preference model, and a hybrid variant with per-item parameters, aimed at overcoming potential model misspecification (Section 5.3).
-We evaluate these models and sampling strategies in two realistic image ordering tasks, one where the preference feedback is simulated using a logistic model, and one where the comparisons have been performed by human annotators (Section 6).
-We show that our sampling strategy compares favorably to active preference learning and bandit baselines, and that our hybrid model benefits both from the low variance of contextual models and the low bias of fitting per-item parameters.
-This manifests in faster convergence in-sample, better generalization to new items, and more efficient tuning when continuing learning on new items.
-2 Ordering with active preference learning
+We propose using a contextual logistic preference model to support efficient in-sample ordering and generalization to new items.
+Our analysis yields the first bound on the expected ordering error achievable given a collected set of comparisons (Section 4).
+This result justifies an active sampling principle that accounts for both epistemic and aleatoric uncertainty which we implement in a greedy deterministic algorithm called GURO (Section 5).
+We further propose a hybrid variant of the contextual preference model, compatible with GURO as well as existing sampling strategies, that overcomes model misspecification by adding per-item parameters (Section 5.1).
+We evaluate GURO and baseline algorithms in four diverse ordering tasks, three of which utilize comparisons performed by human annotators (Section 6).
+Our sampling strategy compares favorably to active preference learning baselines, and our hybrid model benefits both GURO and other sampling criteria, achieving the low variance of contextual models and the low bias of fitting per-item parameters.
+This results in faster convergence in-sample, better generalization to new items, and efficient continual learning when new items are added.
+2 Ordering items with active preference learning
 Our goal is to learn an ordering of items $\mathcal{I}$ according to an unobserved score $y_{i}\in\mathbb{R}$, defined for each item $i\in\mathcal{I}$.
-The ground-truth ordering of $\mathcal{I}$ is determined by a comparison function $\pi_{ij}\coloneqq\mathds{1}[y_{i}>y_{j}]$ such that $\pi_{ij}=1$ indicates that $i$ is ranked higher than $j$.
+The ground-truth ordering of $\mathcal{I}$ is determined by a comparison function $\pi_{ij}\coloneqq\mathds{1}[y_{i}>y_{j}]$, where $\pi_{ij}=1$ indicates that $i$ ranks higher than $j$.
 We assume there are no ties.
@@
 $R_{\mathcal{I}}(h)=\frac{2}{n(n-1)}\sum_{i\neq j\in\mathcal{I}}\mathds{1}[h(i,j)\neq\pi_{ij}]~{},$ (1)
-where $n=|\mathcal{I}|$. This error is equivalent to the normalized Kendall’s Tau distance (Kendall, 1948). Our analysis may be readily generalized to non-uniform distributions.
-We learn $h$ from preference feedback—noisy pairwise comparisons $C_{ij}\in\{0,1\}$ for items $(i,j)$ related to their score, for example, provided by one or more human annotators.
-$C_{ij}=1$ indicates that the annotator perceived that item $i$ has a higher score than $j$, i.e., that they prefer $i$ over $j$.
+where $n=|\mathcal{I}|$. This error is equivalent to the normalized Kendall’s Tau distance (Kendall, 1948).
+Hypotheses $h$ are learned from preference feedback—noisy pairwise comparisons $C_{ij}\in\{0,1\}$ for items $(i,j)$ related to their score, for example, provided by human annotators.
+$C_{ij}=1$ indicates that an annotator perceived that item $i$ has a higher score than $j$, i.e., that they prefer $i$ over $j$.
 Our goal is to minimize the ordering error $R_{\mathcal{I}}(h)$ for a fixed budget $T\geq 1$ of adaptively chosen comparisons.
-We are particularly interested in algorithms that are guaranteed to generalize out-of-sample to unseen items.
-In this setting, only a subset of $\mathcal{I}_{D}\subseteq\mathcal{I}$ is available during learning, and the learner must generalize to $\mathcal{I}\setminus\mathcal{I}_{D}$.
+We are interested in contextual problems, where each item $i\in\mathcal{I}$ is endowed with item-specific attributes $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$.
+As we will see, this permits more sample-efficient ordering and learning algorithms that can order items out-of-sample, trained on comparisons of a subset of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and generalizing to $\mathcal{I}\setminus\mathcal{I}_{D}$.
 Ordering algorithms based only on preference feedback cannot solve this problem since observed comparisons are uninformative of new items.
-To overcome this, we let each item $i\in\mathcal{I}$ be endowed with item-specific attributes $x_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$.
-The active preference learning scenario proceeds as follows:
-  1. 1.
-    A learner is given an annotation budget $T$, a pool of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and per-item attributes $x_{i}$ for $i\in\mathcal{I}_{D}$.
-  2. 2.
-    Over rounds $t=1,...,T$, the learner requests a comparison of two items $i_{t},j_{t}\in\mathcal{I}_{D}$ according to a sampling criterion and receives noisy binary preference feedback $c_{t}\sim p(C_{ij})$, independently of previous comparisons.
-  3. 3.
-    After $T$ rounds, the learner returns an estimated comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$.
-We denote the history of accumulated observations (pairs of items and comparisons) up until and including time $t$ by $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$.
-To learn a good model by time $T$, an active learner must use all information in $D_{t}$ to inform its sampling criterion.
-A natural baseline is the uniform criterion, which selects any pair $i\neq~{}j\in\mathcal{I}_{D}$ with equal probability, ignorant of $D_{t}$.
-Uniform sampling gives equal weight to item pairs with similar scores and pairs with vastly different scores, which may result in many comparisons having minimal impact on the final estimate $h$.
-Logistic preference feedback.
-We analyze the case where comparisons $C_{ij}$ are generated according to a logistic model applied to the difference between item scores, $p(C_{ij}=1)=\sigma(y_{i}-y_{j})~{},$ where $\sigma(z)=1/(1+e^{-z})$ is the logistic function and $\dot{\sigma}(z)$ its derivative at $z$.
-Note that $\dot{\sigma}(z)$ is also the variance of a Bernoulli variable parameterized by $\sigma(z)$.
-We are particularly interested in the case where $y_{i}$ is a linear function of item attributes, $y_{i}=\theta_{*}^{\top}x_{i}~{},$ with $\theta_{*}\in\mathbb{R}^{d}$ the ground-truth coefficients.
-In this case, comparisons are determined by a logistic regression model applied to the difference vector between item attributes $z_{ij}\coloneqq x_{i}-x_{j}$,
+Our active preference learning scenario proceeds as follows: 1) A learner is given an annotation budget $T$, a pool of items $\mathcal{I}_{D}\subseteq\mathcal{I}$ and item attributes $x_{i}$ for $i\in\mathcal{I}_{D}$.
+2) Over rounds $t=1,...,T$, the learner requests a comparison of two items $i_{t},j_{t}\in\mathcal{I}_{D}$ according to a sampling criterion and receives noisy binary preference feedback $c_{t}\sim p(C_{ij})$, independently of previous comparisons.
+3) After $T$ rounds, the learner returns a comparison function $h:\mathcal{I}\times\mathcal{I}\rightarrow\{0,1\}$.
+We denote the history of accumulated observations until and including time $t$ by $D_{t}=((i_{1},j_{1},c_{1}),...,(i_{t},j_{t},c_{t}))$.
+We assume that comparisons $C_{ij}$ follow a logistic model applied to the difference between item scores, $p(C_{ij}=1)=\sigma(y_{i}-y_{j}),$ the so-called Bradley-Terry model (Bradley and Terry, 1952), which assumes linear stochastic transitivity (Oliveira et al., 2018).
+Throughout, $\sigma(z)=1/(1+e^{-z})$ and $\dot{\sigma}(z)$ its derivative at $z$.
+Specifically, we study the case where $y_{i}$ is a linear function of item attributes, $y_{i}=\theta_{*}^{\top}x_{i}~{},$ with $\theta_{*}\in\mathbb{R}^{d}$ the ground-truth coefficients.
+Thus, comparisons are determined by a logistic regression model applied to the attribute difference vector $z_{ij}\coloneqq x_{i}-x_{j}$,
 $p(C_{ij}=1)=\sigma(\theta_{*}^{\top}z_{ij})~{}.$ (2)
-This model is a special case of the family of parametric models considered in Heckel et al.
-(2018), and is equivalent to the logistic model often used in learning to rank (Burges et al., 2005).
-It satisfies linear (and thus strong) stochastic transitivity (Oliveira et al., 2018).
-A similar score function is assumed also in Chu and Ghahramani (2005).
+We face two kinds of uncertainty when actively learning the model in (2): epistemic and aleatoric.
+Epistemic uncertainty, or model uncertainty, is the uncertainty about the true parameter $\theta_{*}$, while aleatoric uncertainty is the irreducible uncertainty about labels due to noisy annotation.
 3 Related work
-Preference learning (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005) is related to the problem of learning to rank (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012).
-Solving preference learning using as few, adaptively chosen comparisons as possible may be posed as an active learning (Brinker, 2004; Long et al., 2010; Silva et al., 2014; Ling et al., 2020) or bandit problem.
-Non-contextual active learners, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018), and Probe-Rank (Lou et al., 2022) produce in-sample orderings, but must be updated if new items are to be ranked.
-This is mitigated by contextual algorithms that exploit structure between items, such as BALD (Houlsby et al., 2011).
-BALD, when applied to preference learning, samples items by maximizing the expected information gain, a commonly used criterion in Bayesian optimal design (Lindley, 1956).
-Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information.
-Chaudhuri et al.
-(2015) analyze active learning for general maximum likelihood estimation, and Ash et al.
-(2021) extend this idea for models parameterized by deep neural networks.
-Bandit algorithms with relative or dueling feedback (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022), have been proposed both in contextual (Dudík et al., 2015) and non-contextual settings (Yue et al., 2012).
+Active Preference Learning:
+Preference learning (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005) is related to the problem of learning to rank (Burges et al., 2005; Busse et al., 2012).
+When using adaptively chosen comparisons it may be posed as an active learning or bandit problem (Brinker, 2004; Long et al., 2010; Silva et al., 2014; Ling et al., 2020).
+Non-contextual active learners, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018), and Probe-Rank (Lou et al., 2022) produce in-sample preference orderings, but must be updated if new items are to be ranked.
+Contextual algorithms, such as BALD (Houlsby et al., 2011), mitigate this by exploiting item structure and Kirsch and Gal (2022) show that many recently proposed contextual active learning strategies may be unified in a framework based on Fisher information.
+Similarly, methods have been proposed to recover a linear preference model by adaptively sampling paired comparisons (Qian et al., 2015; Massimino and Davenport, 2021; Canal et al., 2019).
+Still, this setting differs from ours in that we emphasize recovering the full ordering, not perfectly estimating the parameters.
+While it is true that knowing the parameters is sufficient to order the list, reducing uncertainty for all parameters equally will likely be wasteful (see Section 4).
+Ailon (2011) offer guarantees for ordering using contextual features in the noiseless setting, while Jamieson and Nowak (2011) analyze the setting where noise is unrelated to item similarity.
+Bandits:
+Bandit algorithms with relative or dueling feedback (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022) also learn from pairwise comparisons, and have been proposed both in contextual (Dudík et al., 2015) and non-contextual settings (Yue et al., 2012) to minimize regret or identify top-$k$ items.
 Bengs et al.
-(2022) proposed CoLSTIM, a contextual dueling bandit for regret minimization under linear stochastic transitivity, matching the model in (2).
-Di et al.
+(2022) proposed CoLSTIM, a contextual dueling bandit for regret minimization under linear stochastic transitivity, matching (2), and Di et al.
 (2023) gave variance-aware regret bounds for this setting.
-However, as argued in Appendix E, algorithms that identify the top-$k$ items as fast as possible, such as pure exploration bandits (Fang, 2022; Jun et al., 2021), can be arbitrarily bad at learning a full ordering.
-Related are also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits, and Wu et al.
-(2023b) who minimize the Borda regret.
-Our analysis uses techniques from generalized linear bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020).
-For an extended review of related work, see Appendix B.
-4 A bound on the ordering error
-We give an upper bound on the ordering error $R_{\mathcal{I}}(h)$ for a hypothesis $h$ fit using a fixed budget of comparisons, $T$.
+However, algorithms that find the top-$k$ items, such as pure exploration bandits (Fang, 2022; Jun et al., 2021), can be arbitrarily bad at learning a full ordering (see Appendix D).
+Related are also George and Dimitrakakis (2023) who learn Kemeny rankings in non-contextual dueling bandits, and Wu et al.
+(2023b) who minimize Borda regret.
+Zhu et al.
+(2023) studies the problem of estimating a preference model from offline data.
+Our analysis uses techniques from logistic bandits (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020).
+RLHF:
+Preference learning is commonly used when training large language models through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Wu et al., 2023a).
+In this line of work, Zhu et al.
+(2023) provide guarantees on the sample complexity of learning a preference model from offline data.
+They leverage similar tools from statistical learning and bandits as we do.
+In contrast to their work, we provide sampling strategies for the online setting.
+Mehta et al.
+(2023) consider active learning for RLHF in a dueling bandit framework where the goal is to optimize a contextual version of the Borda regret.
+Concurrent work by Mukherjee et al.
+(2024) and Das et al.
+(2024) studies a similar problem, as we do here, in the RLHF setting but with the objective to identify an optimal policy in a contextual bandit with dueling feedback.
+In contrast to their objective, we are interested in recovering the ordering of items.
+Das et al.
+(2024) use similar bandit techniques as we do, and their selection criterion, when adapted for ordering, corresponds to our NormMin baseline (see Section 6).
+4 Which comparisons result in a good ordering?
+We give an upper bound on the ordering error $R_{\mathcal{I}}(h)$ for a hypothesis $h$ fit using a set of $T$ comparisons.
+The bound is retrospective, attempting to answer the question “if we collect comparisons $D_{T}$, how good is our resulting model at ordering the items in $\mathcal{I}$?”.
+In Section 5, we use insights from the result to design an active learning algorithm.
 We restrict our analysis to the logistic model in (2) and denote by $R(\theta)\equiv R_{\mathcal{I}}(h_{\theta})$ the risk of the hypothesis defined by $h_{\theta}(i,j)=\mathds{1}[\theta^{\top}z_{ij}>0]$.
-Let $\theta_{t}$ be the maximum-likelihood estimate (MLE) of the true parameter $\theta_{*}$, defined in (9) in the Appendix, fit to $t$ rounds of feedback, $D_{t}$.
-Fix $i,j\in\mathcal{I}$ and recall that $z_{ij}=x_{i}-x_{j}$.
-For convenience, we define $z_{t}\equiv z_{i_{t}j_{t}}$ as the difference between attributes for the pair of items selected at round $t$.
-Now, let $\Delta_{ij}>0$ be a lower bound on the margin of comparison such that $\left|\sigma(z_{ij}^{\top}\theta_{*})-1/2\right|>\Delta_{ij}$ for all $i,j$.
-Further, define $\Delta_{*}=\min_{ij}\Delta_{ij}$.
-Let $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$ be the Hessian of the negative log-likelihood of observations at time $t$ under (2), given the parameter $\theta$.
-We define
-$\tilde{\mathbf{H}}_{t}(\theta):=\sum_{i,j}w_{t}(i,j)\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}$
-with $w_{t}(i,j))=N_{t}(i,j)/t$ where $N_{t}(i,j)$ denotes the number of times the pair $(i,j)$ has been queried up to time $t$.
-Hence, $\mathbf{H}_{t}(\theta_{t})=t\tilde{\mathbf{H}}_{t}(\theta_{t})$.
-We further note that $\mathbf{H}_{t}(\theta)$ is the observed Fisher information, and define the norm $||x||_{V}$, for a square matrix $V$, as $\|x\|_{V}=\sqrt{x^{\top}Vx}$.
+Recall that $z_{ij}=x_{i}-x_{j}$ for $i,j\in\mathcal{I}$, and define $z_{t}\equiv z_{i_{t}j_{t}}$ as the difference between attributes for the pair of items selected at round $t$.
+Let $\theta_{t}$ be the maximum-likelihood estimate (MLE) fit to $t$ rounds of feedback, $D_{t}$
+$\theta_{t}=\operatorname*{arg\,max}_{\theta}\sum_{s=1}^{t}\left(c_{s}\log\sigma(\theta^{\top}z_{s})+(1-c_{s})(1-\sigma(\theta^{\top}z_{s}))\right)~{}.$ (3)
+Let $\Delta_{ij}>0$ lower bound the margin of comparison, $|\sigma(z_{ij}^{\top}\theta_{*})-1/2|>\Delta_{ij}$ for all $i,j\in\mathcal{I}$ and define $\Delta_{*}=\min_{i\neq j}\Delta_{ij}/|i-j|$.
+Next, let $\mathbf{H}_{t}(\theta):=\sum_{s=1}^{t}\dot{\sigma}(z_{s}^{\top}\theta)z_{s}z_{s}^{\top}$ be the Hessian of the negative log-likelihood of observations at time $t$ under (2), given the parameter $\theta$, also known as observed Fisher information.
+We define $\tilde{\mathbf{H}}_{t}(\theta):=\frac{1}{t}\mathbf{H}_{t}(\theta)$.
+For a square matrix $V$, we define $\|x\|_{V}=\sqrt{x^{\top}Vx}$.
 We make the following assumptions for our analysis:
@@
 Assumption 3.
-$\mathbf{H}_{T}(\theta_{T})$ and $\mathbf{H}_{T}(\theta_{*})$ have full rank and their minimum eigenvalues are larger than some $\lambda_{0}>0$.
-Assumptions 1 and 2 are required for our analysis but the learner does not have to know these bounds.
-Assumption 3 is naturally satisfied, for sufficiently large $T$, by most reasonable sampling strategies, or can be ensured by allowing for a burn-in phase of $O(d)$ uniformly drawn samples in the beginning of an adaptive strategy.
+$\mathbf{H}_{T}(\theta_{T})$ and $\mathbf{H}_{T}(\theta_{*})$ have full rank and minimum eigenvalues larger than $\lambda_{0}>0$.
+Assumption 1 implies that $\theta_{*}$ lies in some ball with radius $S$ and cannot have unbounded coefficients.
+Assumption 2 states that there exists an upper bound on the norm of the feature vectors.
+This assumption is trivially satisfied whenever we have a finite set of data points.
+Both assumptions 1 and 2 are standard in the bandit literature and only required for our analysis.
+Assumption 3 is naturally satisfied for sufficiently large $T$ by any sampling strategy with support on $d$ linearly independent vectors, or can be ensured by allowing for a burn-in phase of $d$ samples in the beginning of an adaptive strategy.
+Assumption 3 ensures the uniqueness of $\theta_{t}$.
 We start by stating the following concentration result for the deviation of $\sigma(z_{ij}^{\top}\theta_{T})$ from the true probability $\sigma(z_{ij}^{\top}\theta_{*})$.
-The proof of Lemma 1 is found in Appendix D and builds on results for optimistic algorithms in logistic multi-armed bandits (Filippi et al., 2010; Faury et al., 2020).
+The proof of Lemma 1 is found in Appendix C and builds on results for optimistic algorithms in logistic multi-armed bandits (Filippi et al., 2010; Faury et al., 2020).
 Lemma 1 (Concentration Lemma).
 Define, for all pairs of items $i,j\in\mathcal{I}$, and any $\Delta>0$,
-$\alpha_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta^{2}T}{8dC_{1}\left(\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T}))}\right)^{2}}\Bigg{)},\;\;\beta_{ij}(\Delta)$ $:=\exp\Bigg{(}\frac{-\Delta T}{dC_{1}\left(\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)^{2}}\Bigg{)}~{}.$
+$\alpha_{ij}(\Delta):=\exp\Bigg{(}\frac{-\Delta^{2}T}{8dC_{1}(\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})^{2}}\Bigg{)},\;\;\;\beta_{ij}(\Delta):=\exp\Bigg{(}\frac{-\Delta T}{dC_{1}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}}\Bigg{)}~{}.$
 Then, if $\alpha\coloneqq\alpha_{ij}(\Delta),\beta\coloneqq\beta_{ij}(\Delta)$ and $\alpha,\beta\leq\frac{1}{4dT}$,
 $P\left(|\sigma(z_{ij}^{\top}\theta_{*})-\sigma(z_{ij}^{\top}\theta_{T})|>\Delta\right)\leq 2dT\left(\alpha+\beta\right).$
-Here $C_{1}$ is a constant depending on $S,\lambda_{0},Q$ from Assumptions 13 (see Appendix D for definition).
+$C_{1}$ depends on $S,\lambda_{0},Q$ from Assumptions 13 (see Appendix C for definition and proof).
 The concentration result in Lemma 1 is verifiable (given by observables) since the upper bound depends only on the maximum likelihood estimate $\theta_{T}$ at time $T$, not on $\theta_{*}$.
-We present a sharper, unverifiable bound in Appendix D which instead depends on $\theta_{*}$ but does not suffer from the explicit scaling with $d$ in the definitions of $\alpha$ and $\beta$.
-The bound in Lemma 1 can also be expressed in terms of $\mathbf{H}^{-1}_{T}(\theta_{T})$ by using the equality $||z_{ij}||^{2}_{\mathbf{H}^{-1}_{T}(\theta_{T})}=\frac{1}{T}||z_{ij}||^{2}_{\hat{\mathbf{H}}^{-1}_{T}(\theta_{T})}$
-As long as our sampling strategy ensures that the minimum eigenvalue of $\tilde{\mathbf{H}}_{t}(\theta_{t})$ does not tend to zero, i.e., the strategy is strongly consistent (Chen et al., 1999), we have $\alpha_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}^{2}T/(\dot{\sigma}(z_{ij}^{\top}\theta_{T})^{2}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})\right)$ and $\beta_{ij}(\Delta_{ij})\sim\exp\left(-\Delta_{ij}T/\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}\right)$.
+We present a sharper, unverifiable bound in Appendix C which instead depends on $\theta_{*}$ but does not suffer from the explicit scaling with $d$ in the definitions of $\alpha$ and $\beta$.
+The bound in Lemma 1 can also be expressed in terms of $\mathbf{H}^{-1}_{T}(\theta_{T})$ by using the equality $||z_{ij}||^{2}_{\mathbf{H}^{-1}_{T}(\theta_{T})}=\frac{1}{T}||z_{ij}||^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}.$ As long as our sampling strategy ensures that the minimum eigenvalue of $\tilde{\mathbf{H}}_{t}(\theta_{t})$ does not tend to zero, i.e., the strategy is strongly consistent (Chen et al., 1999), we have $\alpha_{ij}(\Delta_{ij})\sim\exp[-\Delta_{ij}^{2}T/(\dot{\sigma}(z_{ij}^{\top}\theta_{T})^{2}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})})]$ and $\beta_{ij}(\Delta_{ij})\sim\exp[-\Delta_{ij}T/\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}]$.
 Since $\Delta^{2}_{ij}<\Delta_{ij}<1/2$ by definition, we can view $\alpha$ as the first-order term and $\beta$ as the second-order term of our bound.
@@
 Lemma 1 also tells us that the hardest pairs to guarantee a correct ordering for are the ones with both high aleatoric uncertainty under the MLE model, e.g., where annotators disagree or labels are noisy, captured by $\dot{\sigma}(z_{ij}^{\top}\theta_{T})$, as well as high epistemic uncertainty captured by $||z_{ij}||_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$.
-A direct consequence of Lemma 1 is the following bound,
+A direct consequence of Lemma 1 is the following bound on the ordering error of $h_{\theta_{T}}$ over $\mathcal{I}$,
 $\mathbb{E}[R(\theta_{T})]\leq\sum_{i\neq j}\frac{2\min\left\{2dT\left(\alpha_{ij}(\Delta_{ij})+\beta_{ij}(\Delta_{ij})\right),1\right\}}{n(n-1)}.$
-The RHS in the above inequality can be bounded further by utilizing that $\Delta_{ij}\geq|i-j|\Delta_{*}$.
-Together with Markov’s inequality, this yields the following bound on $P\left(R(\theta_{T})\geq\epsilon\right)$.
+The right-hand side in the above inequality can be bounded further by utilizing that $\Delta_{ij}\geq|i-j|\Delta_{*}$.
+Together with Markov’s inequality, this yields the following bound on $P(R(\theta_{T})\geq\epsilon)$.
 Theorem 1 (Upper bound on the ordering error).
@@
 where $\alpha_{*}$ and $\beta_{*}$ decay exponentially with $T$.
-A proof is given in Appendix D.
-Theorem 1 suggests that the probability of $R(\theta_{T})\geq\epsilon$ decays exponentially with a rate that depends on the quantities $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$ and $\max_{i,j}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$.
+Theorem 1 suggests that the probability of $R(\theta_{T})\geq\epsilon$ decays exponentially with a rate that depends on the quantities $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$ and $\max_{i,j}\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$.
 Both quantities are random variables that depend on the particular sampling strategy that yields $\mathbf{H}_{t}$.
-Focusing on the leading term, $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{*})}$, Theorem 1 suggests that an active learner should gather data to minimize this quantity and obtain the smallest possible bound.
-The factor $\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$, is related to active sampling criteria based on Fisher information (Kirsch and Gal, 2022); it is the weighted norm of $z_{ij}$ w.r.t.
-the inverse of the observed Fisher information.
+Focusing on the leading term, $\max_{i,j}\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\tilde{\mathbf{H}}^{-1}_{t}(\theta_{T})}$, Theorem 1 suggests that an active learner should gather data to minimize this quantity and obtain the smallest possible bound.
+The factor $\|z_{ij}\|^{2}_{\tilde{\mathbf{H}}^{-1}_{T}(\theta_{T})}$ is the weighted norm of $z_{ij}$ w.r.t.
+the inverse of the observed Fisher information (cf.
+Kirsch and Gal (2022)).
 It controls the shape of the confidence ellipsoid around $\theta_{T}$ and the width of the confidence interval around $\theta_{T}^{\top}z_{ij}$.
 The leading term in Theorem 1 re-scales this quantity with aleatoric noise under the MLE estimate $\theta_{T}$.
-Regularized maximum-likelihood.
-Throughout this section, we have assumed that $\theta_{T}$ is the maximum likelihood estimate and that $\mathbf{H}(\theta_{T})$ has full rank.
-This can be relaxed by considering $\ell_{2}$ (Ridge) regularization where $\theta_{\lambda_{0},T}$ is the optimum of the regularized log-likelihood with regularization $\lambda_{0}\mathbf{I}$ and $\mathbf{H}_{\lambda_{0}}(\theta_{\lambda_{0},T})=\sum_{s=1}^{T}\dot{\sigma}(z_{s}^{\top}\theta_{\lambda_{0},T})z_{s}z_{s}^{\top}+\lambda_{0}\mathbf{I}$.
-The same machinery used to prove Lemma 1 (Filippi et al., 2010; Faury et al., 2020) can be applied to this regularized version and result in small changes to the final bound.
-Other link functions.
-We have focused on the logistic regression model as it is often used to model preferences.
-It is possible to derive similar results for other generalized linear models, with other link functions, $\mu(z_{ij}^{\top}\theta_{*})$.
-For this, we could use the inequality $\mathbf{H}(\theta)\geq\kappa^{-1}\mathbf{V}$ with $\mathbf{V}=\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ and $\kappa\geq 1/\min_{z_{ij}}\dot{\mu}(z_{ij}^{\top}\theta_{*})$.
-To the best of our knowledge, this would yield a scaling of $\sim\exp(-\Delta^{2}T/\kappa)$ where, unfortunately, $\kappa$ might be very large.
-For discussion on the dependence on $\kappa$ in generalized linear bandits, see Lattimore and Szepesvári (2020, Chapter 19).
-5 Algorithms
-We present two active preference learning algorithms for ordering a list of items: GURO, a deterministic algorithm based on greedy minimization of the bound in Theorem 1, and BayesGURO, a stochastic version where we sample $\theta$ from a posterior distribution.
-Both are presented in Algorithm 1.
-As we will see, they both lead to similar sampling criteria.
-We begin by studying preference models in the form $\sigma(\theta^{\top}z_{ij})$ and return in Section 5.3 to variations on the model parameterization.
-Algorithm 1 Greedy Uncertainty Reduction for Ordering
+This suggests that higher epistemic (model) certainty is needed in directions with high aleatoric uncertainty—where item similarity increases noise in comparisons.
+In Appendix C.3, we comment on i) generalizations to regularized preference models, ii) applications to generalized linear models with other link functions, iii) lower bounds on the ordering error, and iv) an algorithm-specific upper bound.
+5 Greedy uncertainty reduction for ordering (GURO)
+We present an active preference learning algorithm based on greedy minimization of the bound in Theorem 1, called GURO.
+We begin with fully contextual preference models of the form $\sigma(\theta^{\top}z_{ij})$ and return in Section 5.1 to parameterization variants to reduce the effects of model misspecification.
+Algorithm 1 Greedy Uncertainty Reduction for Ordering (GURO), [BayesGURO]
 1:Training items $\mathcal{I}_{D}$, attributes ${\bf X}=\{x_{i}\}_{i\in\mathcal{I}_{d}}$
@@
 3:for $t=1,...,T$ do
-4:     Draw $(i_{t},j_{t})$ according to (4) or (5)
+4:     Draw $(i_{t},j_{t})$ based on $\theta_{t}$ according to (5) [or (7)]
 5:     Observe $c_{t}$ from noisy comparison (annotator)
 6:     $D_{t}=D_{t-1}\cup\{i_{t},j_{t},c_{t})\}$
-7:     if GURO then
-8:          Let $\theta_{t}=\mathrm{MLE}(D_{t})$,    (9) in the Appendix
-9:     else if BayesGURO then
-10:          Let $\theta_{t}=\mathrm{MAP}(D_{t})$,    (10) in the Appendix
-11:     end if
-12:end for
-13:Return $h_{T}$
-5.1 Greedy worst-pair uncertainty reduction
-The main component of the bound in Theorem 1 to be controlled by an active learning algorithm is
+7:     $\theta_{t}=\mathrm{MLE}(D_{t})$ according to (3) [or $\theta_{t}=\mathrm{MAP}(D_{t})$ as in (11) in the Appendix]
+8:end for
+9:Return $h_{T}$
+The main component of the bound in Theorem 1 to be controlled by an active learner is the term
 $\max_{i,j\in\mathcal{I}}\;\dot{\sigma}(z_{ij}^{\top}\theta_{T})\|z_{ij}\|_{\mathbf{H}^{-1}_{T}(\theta_{T})}~{},$ (4)
-which represents the highest uncertainty in the comparison of any items $i,j\in\mathcal{I}$ under the model $\theta_{T}$. A smaller value of (3) yields a smaller bound and a stronger guarantee.
+which represents the highest uncertainty in the comparison of any items $i,j\in\mathcal{I}$ under the model $\theta_{T}$.
+A smaller value of (4) yields a smaller bound and a stronger guarantee.
 Recall that, for any $t=1,...,T$, $\theta_{t}$ is the MLE estimate of the ground-truth parameter $\theta_{*}$ with respect to the observed history $D_{t}$.
 Both factors in (4) are determined by the sampling strategy that yielded the item pairs $(i_{t},j_{t})$ in $D_{T}$ and, therefore, $\mathbf{H}_{T}$ and $\theta_{T}$ (the results of comparisons $c_{ij}$ are outside the control of the algorithm, but $z_{ij}$ are known).
-Direct minimization of (3) is not feasible without access to comparisons $c_{ij}$ and their likelihood under $\theta_{T}$.
+Direct minimization of (4), for a subset $\mathcal{I}_{D}$, is not feasible without access to comparisons $c_{ij}$ and their likelihood under $\theta_{T}$.
 Instead, we adopt a greedy, alternating approach: In each round, a) a single pair is sampled for comparison by maximizing (4) under the current model estimate, and b) $\theta_{t}$ is recomputed based on $D_{t}$.
@@
 $i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i\neq j}\dot{\sigma}(z_{ij}^{\top}\theta_{t-1})\|z_{ij}\|_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}~{}.$ (5)
-We refer to this criterion as GURO for Greedy Uncertainty Reduction for Ordering, since sampling in this way reduces the uncertainty of $\theta_{t}$ in the direction of $z_{ij}$.
+We refer to this sampling criterion as Greedy Uncertainty Reduction for Ordering (GURO), since it reduces the uncertainty of $\theta_{t}$ in the direction of $z_{ij}$.
 To see this, consider the change of $\mathbf{H}_{t}(\theta_{t})$ after a single play of $i_{t},j_{t}$.
-The Sherman-Morrison formula (Sherman and Morrison, 1950) yields, with $z_{t}\coloneqq z_{i_{t}j_{t}}$,
-$\mathbf{H}_{t}^{-1}(\theta_{t-1})$ $=\mathbf{H}_{t-1}^{-1}(\theta_{t-1})-\underbrace{\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\frac{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})z_{t}z_{t}^{\top}\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}{1+\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\|z_{t}\|^{2}_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}}}_{=:{\color[rgb]{1,0,0}\xi}}$
-and thus, for all $i<j\in\mathcal{I}$, with $\mathbf{H}_{t-1}=\mathbf{H}_{t-1}(\theta_{t-1})$, it holds that $||z_{ij}||^{2}_{\mathbf{H}_{t}^{-1}(\theta_{t-1})}=||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}-||z_{ij}||^{2}_{{\color[rgb]{1,0,0}\xi}}\leq||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}~{}.$ The inequality is strict for the pair $i_{t},j_{t}$ in (4).
+The Sherman-Morrison formula (Sherman and Morrison, 1950) yields,
+$\mathbf{H}_{t}^{-1}(\theta_{t-1})=\mathbf{H}_{t-1}^{-1}(\theta_{t-1})-\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\frac{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})z_{t}z_{t}^{\top}\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}{1+\dot{\sigma}(z_{t}^{\top}\theta_{t-1})\|z_{t}\|^{2}_{\mathbf{H}_{t-1}^{-1}(\theta_{t-1})}}~{},$ (6)
+where $z_{t}\coloneqq z_{i_{t}j_{t}}$.
+With $\xi$ as the second term in (6), it holds for all $i<j\in\mathcal{I}$, with $\mathbf{H}_{t-1}=\mathbf{H}_{t-1}(\theta_{t-1})$, that $||z_{ij}||^{2}_{\mathbf{H}_{t}^{-1}(\theta_{t-1})}=||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}-||z_{ij}||^{2}_{\xi}\leq||z_{ij}||^{2}_{\mathbf{H}_{t-1}^{-1}}~{}.$ The inequality is strict for the pair $i_{t},j_{t}$ in (5).
 As $\theta_{t}$ converges to $\theta_{*}$, this pair becomes representative of the maximizer of (4) provided there is no major systematic discrepancy between $\mathcal{I}_{D}$ and $\mathcal{I}$.
-A drawback of a deterministic greedy criterion like (3) is that it cannot readily be applied in batched settings since it will sample the same pair of items throughout the batch.
-Next, we offer a Bayesian stochastic alternative.
-5.2 Stochastic Bayesian uncertainty reduction
+Surprisingly, GURO can also be justified from a Bayesian analysis.
 Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$.
-In our second algorithm, we sample items $i_{t},j_{t}$ based on a stochastic estimate of the posterior variance of predictions for their comparison,
-$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i<j}\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]~{}.$ (5)
-Here, $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ is an estimate of the variance of probabilities $\sigma(\theta^{T}z_{ij})$, computed from finite samples drawn from the posterior of $\theta$.
-Estimating the variance in this way both i) allows for tractable implementation, and ii) induces controlled stochasticity in the selection of item pairs.
-We refer to the sampling criterion in (5) as BayesGURO.
-For the logistic model considered in this section, using Laplace approximation with a Normal prior $\mathcal{N}(0,\mathbf{H}^{-1}_{B,0})$ on $\theta$, the Bayesian criterion in (5) is related to the criterion in (4) through the first-order Taylor expansion of the variance:
-$\mathbb{V}_{\theta\mid D_{t}}(\sigma(\theta^{\top}z_{ij}))$ $\approx(\dot{\sigma}(\mathbb{E}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]))^{2}\mathbb{V}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]$
-$=(\dot{\sigma}(\theta_{B,t}^{\top}z_{ij})||z_{ij}||_{\mathbf{H}^{-1}_{B,t}(\theta_{B,t})})^{2}~{},$
-where $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ and $\mathbf{H}_{B,t}$ is the Hessian adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$ (further described in Appendix C.2).
-Thus, to a first-order approximation, for a large number of posterior samples, the two active learning criteria are equivalent, save for the influence of the prior.
-In practice, we find that the Bayesian variant lends itself well to sequential updates of the posterior.
-The choice of prior $p(\theta)$, which could be useful under strong domain knowledge, and the stochasticity of using few posterior samples to approximate $\mathbb{V}$ make the two criteria distinct.
-5.3 Preference models
-We consider logistic preference models aimed at mitigating different sources of error.
-In all variants, we let $h(i,j)=\mathds{1}[f(i,j)>0]$ for some function $f$ approximating the difference in score between items $i$ and $j$.
-Throughout, we estimate $f$ using probabilistic modeling, with a logistic likelihood function $\sigma(f(i,j))\approx p(C_{ij}=1)$, matching (2).
-Our default model (see (2) and Section 4) is fully contextual,
-$f_{\theta}(x_{i},x_{j})=\theta^{\top}(x_{i}-x_{j})~{}.$ (6)
-A strength of this model is that the variance in its estimates grows with $d$, but not with $n=|\mathcal{I}|$.
-As we see empirically (e.g., in Figure 4), this results in much quicker convergence than non-contextual methods.
-This fully contextual model also generalizes to unseen items in $\mathcal{I}$ under modest assumptions on how $\mathcal{I}_{D}$ was sampled.
-The model’s main drawback is two other sources of error: model misspecification bias (error due to the chosen functional form), and noise (error due to insufficiency of $X$ in predicting $C$).
-The former can be mitigated by applying the linear model to a representation function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$,
-$f_{\theta}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j}))~{}.$
-A good representation $\phi$, e.g., from a foundation model, can mitigate misspecification bias, and enable using different data modalities of input.
+A natural active learning strategy is to sample items $i_{t},j_{t}$ for which the model preference is highly uncertain under the posterior distribution,
+$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i<j}\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]~{},$ (7)
+where $\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]$ is a finite-sample estimate of the variance in predictions, computed by sampling from the posterior.
+In Appendix B.3, we show that the first-order Taylor expansion of the true variance is equal to the GURO criterion.
+Hence, we refer to sampling according to (7) as BayesGURO.
+Unlike GURO, BayesGURO can incorporate prior knowledge through $p(\theta)$ and benefits from controlled stochasticity through the empirical estimate $\hat{\mathbb{V}}$, which makes it appropriate for batched algorithms—a deterministic criterion would construct batches of a single item pair.
+Both GURO and BayesGURO are presented in Algorithm 1.
+Computational Complexity:
+Running the algorithms requires $O(n^{2})$ operations each iteration to evaluate the sampling criteria (Equation 5 or 7) on all possible pairs, a problem shared by many active preference learning algorithms (Qian et al., 2015; Canal et al., 2019; Houlsby et al., 2011).
+A way of mitigating this computational complexity is to, at each time step, sample a fixed number of comparisons and only evaluate on these, similar to the approach taken in Canal et al.
+(2019).
+When only looking at a sample of $m\ll n^{2}$ pairs, the complexity is reduced to $O(m)$.
+While making $m$ too small can hurt the sample complexity, we describe in Appendix E how we implemented this sub-sampling strategy to speed up computations in one of our experiments and observed no noticeable change in performance.
+Lastly, we want to highlight that in many realistic scenarios, the computational burden pales in comparison to the time it takes to query an annotator.
+5.1 Preference models for in- and out-of-sample ordering
+Our default preference model $h(i,j)=\mathds{1}[f(i,j)>0]$ is based on a fully contextual scoring function
+$f_{\theta}(x_{i},x_{j})=\theta^{\top}(x_{i}-x_{j})~{},$ (8)
+fit with a logistic likelihood $\sigma(f(i,j))\approx p(C_{ij}=1)$.
+The model’s strength is that the variance in its estimates grows with $d$, but not with $n=|\mathcal{I}|$, often resulting in quicker convergence than non-contextual methods for moderate dimension $d$ (see, e.g., Figure 6).
+The fully contextual model also generalizes to unseen items as long as the attributes for $\mathcal{I}_{D}$ span attributes observed for $\mathcal{I}$.
+The limitations of a fully contextual model are model misspecification (error due to the functional form), and noise (error due to $C$ not being fully determined by $X$).
+The former can be mitigated by applying the linear model to a representation function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$, $f_{\theta}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j})).$ A good representation $\phi$, e.g., from a foundation model, can mitigate misspecification and admit different input modalities.
 As demonstrated in Figure 9 in the Appendix, even a representation pre-trained for a different task can perform much better than a random initialization.
-[^1] It is also feasible to update representations during exploration (Xu et al., 2022), but we do not consider that here.
-Noise due to insufficiencies in $X$ cannot be mitigated by a representation $\phi(x)$; If $X$ does not sufficiently cover aspects considered by annotators, even the Bayes classifier $\mathds{1}[p(C_{ij}\mid X_{i},X_{j})>0.5]$ will have a large ordering error.
-However, for in-sample ordering, we can introduce per-item parameters $\zeta_{i}\in\mathbb{R}$ to the score function, one for each item $i\in\mathcal{I}_{D}$, that compensate for both model misspecification and for uninformative attributes,
+[^1] It is feasible to update representations during exploration (Xu et al., 2022; Singh and Chakraborty, 2021), but we do not consider that here.
+Noise due to insufficiencies in $X$ cannot be mitigated by a representation $\phi(x)$; If annotators consistently compare items based on features $U$ not included in $X$, no function $h(X_{i},X_{j})$ can perfectly order the items.
+However, for in-sample ordering of $\mathcal{I}_{D}$, adding per-item parameters $\zeta_{i}\in\mathbb{R}$ to the scoring function, one for each item $i\in\mathcal{I}_{D}$, can mitigate both misspecification and noise,
 $f_{\theta,\boldsymbol{\zeta}}(x_{i},x_{j})=\theta^{\top}(\phi(x_{i})-\phi(x_{j}))+(\zeta_{i}-\zeta_{j})~{}.$ (9)
-Here, $\zeta_{i}-\zeta_{j}$ corrects the residual of the fully contextual model, which can be assumed small if a) the context captures the most relevant information about the ordering, and b) the functional form $\theta^{\top}\phi(x_{i})$ is nearly well-specified.
-$\zeta_{i}-\zeta_{j}$ alone are sufficient to learn a perfect in-sample ordering, at the cost of increased variance (the full dimension is now $n$ instead of $d$) and poor generalization (since $\zeta_{i}$ are unknown for items $i\not\in\mathcal{I}_{D}$).
-In experiments, the hybrid model in (7) exhibits the best of both contextual and per-item parameterizations: When $\phi$ is poor, the model recovers and competes with non-contextual models (Figure 7); when $\phi$ is good, convergence is as quick as fully contextual models.
+We call this a hybrid model and apply it in “GURO Hybrid” and baselines in experiments.
+The term $\zeta_{i}-\zeta_{j}$ can correct the residual of the fully contextual model, which is small if a) the context captures the most relevant information about the ordering, and b) the functional form $\theta^{\top}\phi(x_{i})$ is nearly well-specified.
+Using $\zeta_{i}-\zeta_{j}$ alone is sufficient in-sample, but has high variance (the dimension is $n$ instead of $d$) and poor generalization ($\zeta_{i}$ are unknown for items $i\not\in\mathcal{I}_{D}$).
+In practice, we use L2 regularization to prevent the model from learning an arbitrary $\theta$ by using the full expressivity of $\zeta_{i}$ (see Appendix E for details).
+Empirically, our hybrid models exhibit the best of both worlds: When $\phi$ is poor, the model recovers and competes with non-contextual models (Figure 9); when $\phi$ is good, convergence matches fully contextual models (Figure 6).
 6 Experiments
-We evaluate the GURO variants in two image ordering tasks, one with logistic (synthetic) preference feedback, and one based on real-world feedback from human annotators.
-We provide a synthetic experiment in Appendix F.3 that includes empirical estimates of the bound in Theorem 1.
-Baselines
-We select diverse baseline algorithms for active sampling of comparisons: BALD (Houlsby et al., 2011), CoLSTIM (Bengs et al., 2022), TrueSkill (Graepel, 2012; Hees et al., 2016), Uniform sampling, and a version of GURO that ignores aleatoric uncertainty (referred to as NormMin).
-Bayesian Active Learning by Disagreements (BALD) is, a priori, considered the strongest baseline since it is a contextual active learning algorithm, unlike the others.
-When applied to preference learning, BALD determines the next comparison by maximizing the following criterion
-$\operatorname*{arg\,max}_{i\neq j\in\mathcal{I}_{D}}H[C_{ij}\mid z_{ij},D_{t}]-\underset{p(\theta\mid D_{t})}{\mathbb{E}}[H[C_{ij}\mid z_{ij},\theta_{t}]]~{},$ (8)
-where $H$ is the conditional entropy.
-The criterion aims to greedily maximize the decrease in posterior entropy, which amounts to reducing the epistemic uncertainty.
-The second term in (8) will downplay the influence of the aleatoric uncertainty.
-This is not always beneficial, as suggested by our results in Section 4, since learners may require several comparisons of high-uncertainty pairs to get the order right.
-CoLSTIM was developed for regret minimization and is not expected to perform well here, but is included to illustrate the mismatch between regret minimization and our setting.
-Uniform sampling is oblivious to the observed data and represents a non-active learner, but remains a competitive baseline in active learning tasks (Yang and Loog, 2018).
-TrueSkill (Herbrich et al., 2006) is a non-contextual skill-rating system that models each item as a Gaussian distribution without leveraging item attributes.
+We evaluate GURO (Algorithm 1) and GURO Hybrid (see Section 5.1) in four image ordering tasks, one with logistic (synthetic) preference feedback, and three tasks based on real-world feedback from human annotators [^2] Our code is available at: https://github.com/Healthy-AI/GURO.
+We provide a synthetic experiment in Appendix E.2 that includes empirical estimates of the bound in Theorem 1.
+The experiments include five diverse baseline algorithms, described next.
+BALD (Houlsby et al., 2011) is a priori the strongest baseline since it is a contextual active learning algorithm, unlike the others.
+Its selection criterion greedily maximizes the decrease in posterior entropy, which amounts to reducing the epistemic uncertainty and includes a term to downplay the influence of aleatoric uncertainty.
+This is not always beneficial, as suggested by our analysis in Section 4, since learners may require several comparisons of high-uncertainty pairs to get the order right.
+CoLSTIM (Bengs et al., 2022) is a contextual bandit algorithm, developed for regret minimization and is not expected to perform well here.
+It is included to illustrate the mismatch between regret minimization and our setting.
+TrueSkill (Herbrich et al., 2006; Graepel, 2012) is a non-contextual skill-rating system that models the score of each item as a Gaussian distribution, disregarding item attributes, and has been adopted in various works to score items based on subjective pairwise comparisons (Larkin et al., 2022; Naik et al., 2014; Sartori et al., 2015).
 We use the sampling rule from Hees et al.
 (2016), designed for ordering.
-Finally, to illustrate the importance of accounting for aleatoric uncertainty, we use a version of GURO called NormMin that ignores the $\dot{\sigma}(z_{ij}^{\top}\theta_{t})$ term and plays the pair maximizing $\|z_{ij}\|_{\mathbf{H}_{t}^{-1}(\theta_{t})}$, i.e., it minimizes the second-order term in Lemma 1.
-See Appendix F.3 for a detailed comparison.
-(a) Kendall’s tau distance between estimated and true orderings.
-(b) Mean generalization error with the 95% confidence interval.
+Finally, we include Uniform sampling, and to illustrate the importance of accounting for aleatoric uncertainty, we use a version of GURO called NormMin that ignores the $\dot{\sigma}(z_{ij}^{\top}\theta_{t})$ term and plays the pair maximizing $\|z_{ij}\|_{\mathbf{H}_{t}^{-1}(\theta_{t})}$, i.e., it minimizes the second-order term in Lemma 1.
+NormMin corresponds to the selection criterion in the concurrent work Das et al.
+(2024), adapted to our problem of finding the correct ordering.
+We refer the reader to Appendix E.2 for a detailed comparison where NormMin performs significantly worse than Uniform on certain problem instances, and Appendix E for details regarding the implementation and the choice of hyperparameters for GURO, BayesGURO, and baselines.
+6.1 Ordering X-ray images under the logistic model
+Our first task (X-RayAge) is to order X-ray images based on perceived age (Ieki et al., 2022) where the preference feedback follows a (well-specified) logistic model.
+We base this experiment on the data from the Kaggle competition ”X-ray Age Prediction Challenge” (Felipe Kitamura, 2023) which contains more than 10 0001000010\ 00010 000 de-identified chest X-rays, along with the person’s true age.
+Features were extracted using the 121-layer DenseNet in the TorchXrayVision package (Cohen et al., 2022) followed by PCA projection, resulting in $35$ features.
+A ridge regression model, $\theta_{*}$, was fit to the true age ($R^{2}\approx 0.67$).
+During active learning, feedback is drawn from $p(C_{ij}=1)=\sigma\left(\theta_{*}^{\top}z_{i,j}\cdot\lambda\right)$, where $\lambda$ (set to 0.1) controls the noise level.
+We only include the fully contextual models here since they are well-specified by design, meaning $\mathcal{I}$ can be ordered using only contextual features.
+(a) Mean $R_{I_{D}}$ with 1-sigma error region.
+(b) Mean generalization error (95% CI)
 Figure 1: X-RayAge.
-Performance of active sampling strategies when comparisons are simulated using a logistic regression model according to (2).
+Performance of active sampling strategies when comparisons are simulated using a logistic model according to (2).
 In-sample Kendall’s Tau distance $R_{I_{D}}$ on 200 images (left) and generalization error $R_{I_{E}}-R_{I_{D}}$ for models trained on 150 images and evaluated on 150 images from a different distribution (right).
 All results are averaged over $100$ different random seeds.
-Implementation details
-GURO, CoLSTIM, and Uniform use the logistic regression implementation in Scikit-learn (Pedregosa et al., 2011) with default Ridge regularization ($C=1$).
-The former two update $\theta_{t}$ every iteration using the full history, $D_{t}$.
-The design matrix for CoLSTIM is updated as in Bengs et al.
-(2022): the confidence width $c_{1}$ was chosen to be $\sqrt{d\log(T)}$, and the perturbed values were generated using the standard Gumbel distribution.
-GURO updates $\mathbf{H}^{-1}_{t}$ every iteration using the Sherman-Morrison formula.
-For BayesGURO and BALD, the posterior $p(\theta\mid D_{t})$ is estimated using the Laplace approximation as described in Bishop and Nasrabadi (2006).
-With this approximation, the covariance matrix is the same as the inverse of the Hessian of the log-likelihood.
-For both methods, the priors $\theta_{B,0}=\mathbf{0}^{d}$ and $\mathbf{H}_{B,0}^{-1}=I_{d}$ were used, and sequential updates were performed every iteration.
-The sample criterion for BALD under a logistic model is given in Appendix C.4.1.
-For BayesGURO, 50 posterior samples were used to estimate $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ for every $z_{ij}$.
-The hybrid algorithms follow the same structure, with a small change in sampling (see Appendix F.2), and a constraint that each per-item parameter $\zeta_{i}$ is independent of other parameters.
-This allows for efficient updates of $\mathbf{H}_{B,t}^{-1}$ by using sparsity in the covariance.
-6.1 Ordering X-ray images under the logistic model
-Our first task (X-RayAge) is to order X-ray images based on perceived age (Ieki et al., 2022) where the preference feedback follows a (well-specified) logistic model.
-We base this experiment on publicly available data from the Kaggle competition ”X-ray Age Prediction Challenge” Felipe Kitamura (2023), hosted by the Radiology and Diagnostic Imaging Society of São Paulo (SPR), which contains more than 10 0001000010\ 00010 000 de-identified chest X-rays (one image per subject), along with the person’s true age.
-Features were extracted using the 121-layer DenseNet in the TorchXrayVision package (Cohen et al., 2022) followed by PCA projection, resulting in $35$ features.
-A ridge regression model, $\theta_{*}$, was fit to the true age ($R^{2}\approx 0.67$).
-During the active learning phase, feedback is drawn from $p(C_{ij}=1)=\sigma\left(\theta_{*}^{\top}z_{i,j}\cdot\lambda\right)$ where $\lambda$ (set to 0.1) controls the noise level.
-We only look at the fully contextual models here since they are well-specified by design, $\mathcal{I}$ can be ordered using only contextual features.
-Table 1: X-RayAge.
-Normalized Kendall’s Tau distance for models ordering $n=200$ X-ray images represented by $d=35$ features at different numbers of comparisons $T\in\{\log n,d,n,n\log n\}$, along with the $95\%$ CI.
-$T=\log n$ $d$ $n$ $n\log n$
-Uniform $0.28\pm 0.07$ $0.19\pm 0.04$ $0.12\pm 0.02$ $0.05\pm 0.01$
-BALD $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.02$ $0.07\pm 0.01$
-CoLSTIM $0.28\pm 0.08$ $0.19\pm 0.04$ $0.13\pm 0.02$ $0.09\pm 0.01$
-NormMin $\mathbf{0.27}\pm 0.07$ $0.19\pm 0.04$ $0.15\pm 0.03$ $0.08\pm 0.01$
-GURO $\mathbf{0.27}\pm 0.07$ $\mathbf{0.16}\pm 0.03$ $\mathbf{0.09}\pm 0.01$ $\mathbf{0.04}\pm 0.01$
-BayesGURO $0.32\pm 0.09$ $0.17\pm 0.03$ $0.10\pm 0.01$ $\mathbf{0.04}\pm 0.01$
 In the first setting, we sub-sample 200 X-ray images uniformly at random from the full set.
 A ground-truth ordering of these elements is derived using the learned linear model.
-Figure 1(a) shows the ordering error over 2 000 iterations while Table 1 highlights the error for different budgets, $T$.
+Figure 1(a) shows the ordering error over 2 000 iterations.
 GURO and BayesGURO perform similarly, both better than the baselines.
-We note that BALD starts off converging about as fast as the GURO algorithm, but plateaus, most likely as a result of actively avoiding comparisons with high aleatoric uncertainty.
-Finally, the poor performance of CoLSTIM highlights the discrepancy between best-arm identification and recovering a complete ordering.
-In Table 1, we see that NormMin performs worse than our proposed algorithms and even worse than Uniform, once the budget is large enough.
-This shows the importance of accounting for aleatoric uncertainty and that active sampling to optimize the second-order term in Lemma 1 is not sufficient.
+BALD starts off converging about as fast as GURO, but plateaus, most likely as a result of actively avoiding comparisons with high aleatoric uncertainty—pairs where annotators disagree in their preferences.
+The poor performance of CoLSTIM highlights the discrepancy between regret minimization and recovering a complete ordering.
 In the second setting, we evaluate how well the algorithms generalize to new items.
@@
 The algorithms were then trained to order the list containing the younger subjects, but were simultaneously evaluated on how well they could sort the list containing the older subjects.
-The simulation was run for 2 00020002\ 0002 000 iterations.
 The continuously measured difference in ordering error evaluated on $I_{E}$ and $I_{D}$ are presented in Figure 1(b).
-As expected, all algorithms are worse at ordering the items in $I_{E}$, $R_{I_{E}}-R_{I_{D}}>0$.
-However, GURO and BayesGURO achieve the lowest average difference.
-Together with Figure 1(a), this means that our proposed algorithms achieve the best in-sample and out-of-sample orderings.
+While all algorithms are worse at ordering items in $I_{E}$, GURO and BayesGURO achieve the lowest average difference.
+Together with Figure 1(a), this means that our proposed algorithms achieved the best in-sample and out-of-sample orderings.
+For completeness, the in-sample performance of algorithms in the generalization experiment in Figure 1(b) are included in Appendix E.2.
+6.2 Ordering items with human preference data
+Table 1: Datasets with preference feedback from annotators. Pretrained models are ResNet34 (He et al., 2016), all-mpnet-base-v2 (Reimers and Gurevych, 2019), and FaceNet (Schroff et al., 2015).
+Dataset $n$ $d$comparisons Data type Embedding Model
+ImageClarity $100$ $63$ 27 7302773027\ 73027 730 Image ResNet34 (Imagenet)
+WiscAds $935$ $162$ 9 52895289\ 5289 528 Text all-mpnet-base-v2
+IMDB-WIKI-SbS $6072$ $75$ 110 349110349110\ 349110 349 Image FaceNet (CASIA-Webface)
+Next, we evaluate our algorithm on three publicly available datasets to study the algorithms’ performance when preference feedback comes from human annotators (see Table 1 for an overview, detailed information of datasets in Appendix E.1).
+The datasets are IMDB-WIKI-SbS (Pavlichenko and Ustalov, 2021), where annotators have stated which of two people appear older, ImageClarity (Zhang et al., 2016), where modified versions of the same image have been compared according to the level of distortion, as well as the extended WiscAds dataset (Carlson and Montgomery, 2017), where labels correspond to which political advertisement is perceived as more negative toward an opponent.
+In all datasets, pairs of items were sampled uniformly for annotation.
+For each experiment, we construct a feature vector $\phi(x_{i})\in\mathbb{R}^{d}$ for all $n$ items using a pre-trained embedding model followed by PCA, applied to reduce computational complexity.
+We restrict algorithms to only query pairs for which an annotation exists and remove the annotation from the pool once queried.
+In cases where multiple annotations exist for the same pair, the feedback is chosen randomly among these.
 {subcaptionblock}
-[t]0.48 Refer to caption {subcaptionblock}[t]0.48 Refer to caption
-Figure 2: Kendall’s Tau distance between estimated and true ordering of 6 07260726\ 0726 072 images.
-Figure 3: Ordering a list $\mathcal{I}_{D}$ of 3 00030003\ 0003 000 images.
-After 10 0001000010\ 00010 000 comparisons, the remaining 3 07230723\ 0723 072 images, $\mathcal{I}\setminus\mathcal{I}_{D}$, are added and $R_{\mathcal{I}}$ is plotted.
-Figure 4: PhotoAge.
-The ordering error when comparisons are made by human annotators and no linear model is assumed.
-The convergence on the training images (left) as well as how well the methods generalize to new images (right).
-To avoid cluttering, only the Bayesian algorithms were included as they lend themselves to sequential updates.
-6.2 Ordering IMDB images with human annotations
-Next, we use the public IMDB Age dataset (Pavlichenko and Ustalov, 2021), to study the algorithms’ performance when preference feedback comes from human annotators (PhotoAge).
-The data contains 9 15091509\ 1509 150 images of actors and 250 249250249250\ 249250 249 pairwise comparisons where annotators have stated which of two people appears older.
-No image pairs were rated twice.
-After removing grayscale and low-resolution images, 6 07260726\ 0726 072 images and 110 349110349110\ 349110 349 comparisons remain.
-We extract features from each image using the Inception-ResNet implemented in FaceNet Schroff et al.
-(2015) followed by PCA, resulting in $d=75$ features per image.
-We generate a ground-truth ordering by sorting images by their Borda score, i.e.
-the fraction of comparisons where an actor was considered to be the oldest; the real ages of the actors were not used (see discussion in Appendix F.1).
-We restrict algorithms to only query pairs for which an annotation exists and remove the pair from the pool once queried.
-In the first setting, all 6 07260726\ 0726 072 images were ordered using 50 0005000050\ 00050 000 comparisons.
-We compare the fully contextual (6) and hybrid (7) GURO models, and TrueSkill, a non-contextual algorithm that only uses per-item parameters, included here since the contextual models may be misspecified.
-The results, see Figure 4, highlight the differences between modeling strategies.
-The fully contextual algorithms initially converge quickly, achieving a rough ordering of the items, before plateauing and not making any real improvements.
-This most likely correlates with the richness of the features, meaning that while they are informative enough to roughly order the list, they are not sufficient for retrieving a more granular ordering.
-TrueSkill on the other hand starts by improving the ordering at a much lower pace, but keeps converging more steadily throughout.
-Perhaps most interesting is the result of the hybrid algorithms, which seemingly reap the benefits of both methods, converging quickly like the contextual, while not hitting a plateau.
-In fact, as discussed in Appendix F.3, the hybrid models perform comparably to TrueSkill even when features are completely uninformative.
-One can also observe that while the choice of adaptive strategy has a somewhat modest impact in the fully contextual case, it has a large effect on the hybrid models.
-The second experiment is a few-shot scenario where after some time, additional images are added to the pool of items.
+[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption{subcaptionblock}[t]0.45 Refer to caption
+Figure 2: ImageClarity. $n=100$, $d=63$.
+Figure 3: WiscAds. $n=935$, $d=162$.
+[…]
+[…]
+Figure 6: The empirical error $\hat{R}_{D^{\prime}}(h)$ on a holdout comparison set $D^{\prime}$ when comparisons are made by human annotators.
+The plots are averaged over $100$ (6,6) or 10 (6,6) seeds, and the shaded area represents one standard deviation above and below the mean.
+For every seed, $10\%$ of comparisons were used for the holdout set.
+In (6) we initially order a list $\mathcal{I}_{D}$ of 3 00030003\ 0003 000 images.
+After 10 0001000010\ 00010 000 comparisons the remaining 3 07230723\ 0723 072 images, $\mathcal{I}\setminus\mathcal{I}_{D}$, are added.
+The images in the ImageClarity dataset have been constructed to have an objective ground truth ordering but this is not the case for WiscAds or IMDB-WIKI-SbS.
+As the ground-truth ordering is generally unknown also in real-world applications, we evaluate methods by the error on a held-out set of comparisons $D^{\prime}$, $\hat{R}_{D^{\prime}}(h)=\frac{1}{|D^{\prime}|}\sum_{(i,j,c)\in D^{\prime}}\mathds{1}[h(i,j)\neq c]$.
+This serves as an empirical analog of Kendall’s Tau distance and a minimizer of $\hat{R}_{D^{\prime}}(h)$ will minimize $R_{\mathcal{I}}(h)$ for sufficiently large $D^{\prime}$, but will not converge toward 00 since there is inherent noise in annotations.
+This metric makes no assumptions on the ground truth ordering unlike the alternative approach of fitting an ordering to all available comparisons, see e.g., Maystre and Grossglauser (2017).
+In Appendix E.2, we show results for the latter that highlight the limitations of estimating a “ground-truth” ordering, as well as the similar results when measuring the distance to the objective ground-truth ordering of the ImageClarity dataset.
+The longest trajectory (single seed) for any algorithm took less than 35hrs to complete on one core of an Intel Xeon Gold 6130 CPU and required at most 10 GB of memory.
+In all experiments, we compare fully contextual (8) and hybrid (9) versions of GURO, BALD, and Uniform, as well as TrueSkill.
+The results of each experiment can be seen in Figure 6.
+Figure 6 shows that the ImageClarity dataset is the easiest to order using contextual (non-hybrid) features.
+This is expected, as features relevant to the level of distortion are low-level.
+In this case, the choice of adaptive strategy has a modest impact on the ordering error.
+Figures 6 and 6 highlight the differences between modeling strategies.
+The fully contextual algorithms initially improve rapidly, achieving a rough ordering of the items, before plateauing and not making any real improvements.
+This indicates that the features are informative enough to roughly order the list, but insufficient for retrieving a more granular ordering.
+The non-contextual TrueSkill converges at a much slower pace but keeps improving steadily throughout.
+Perhaps most interesting are the hybrid algorithms, which seemingly reap the benefits of both methods, improving as quickly as the contextual methods, but avoiding the plateau.
+In fact, in Figure 9 in the Appendix we show that the hybrid models perform comparably to TrueSkill even when features are completely uninformative.
+The limitations of BALD are most noticeable in the fully contextual case, where it plateaus at a higher error compared to GURO and Uniform.
+This is however not as prominent when we use BALD in conjunction with our hybrid model, likely a result of the increased dimensionality of the model causing BALD Hybrid to attribute more of the observed errors to model uncertainty.
+While this initially causes the algorithm to avoid fewer comparisons that are subject to aleatoric uncertainty, the final iterations in Figure 6 suggest that BALD Hybrid can still run into this issue given enough samples.
+In all experiments, GURO and GURO Hybrid perform better than or similar to our baselines, never worse.
+Additionally, Figures 6 and 6 showcase how our hybrid model can increase performance when used with existing sampling strategies, such as BALD or Uniform.
+The final experiment, visible in Figure 6, is a few-shot scenario where after some time, additional images are added to the pool of items.
+IMDB-WIKI-SbS was used as it contained the highest number of both images and comparisons.
 The initial pool consists of 3 00030003\ 0003 000 images sampled from the dataset.
 After 10 0001000010\ 00010 000 steps, the remaining 3 07230723\ 0723 072 images were added to the pool.
-Figure 4 shows how the ordering error is affected by the addition of new data for all three types of models.
-We observe that the increase in error of the fully contextual model is very slight, most likely a result of added samples being drawn from the same distribution.
-In the case of TrueSkill, the error increases drastically as a result of the algorithm not having seen these items before and having no way of generalizing the results of previous comparisons to them.
+The results again emphasize the differences between our three types of models; the increase in error of the fully contextual model is very slight, likely a result of added samples being drawn from the same distribution.
+For TrueSkill, the error increases drastically as a result of the algorithm not having seen these items before and having no way of generalizing the results of previous comparisons to them.
 Lastly, the hybrid algorithms seem to be moderately affected.
-The error does increase as the model has not yet tuned any of the added per-item parameters, but one can also observe that the extent of this is much lower than for TrueSkill as the model can provide a rough ranking of the out-of-sample elements using the contextual features.
+The error increases as the model has not yet tuned any of the added per-item parameters, but the extent is much smaller than for TrueSkill as the model can provide a rough ranking of the out-of-sample elements using the contextual features.
 7 Conclusion
-We have demonstrated the benefit of utilizing contextual features in active preference learning to efficiently order a list of items.
+We have demonstrated the benefits of utilizing contextual features in active preference learning to efficiently order a list of items.
 Empirically, this leads to quicker convergence, compared to non-contextual methods, and allows algorithms to generalize out-of-sample.
-We derived an upper bound on the ordering error and used it to design two active sampling strategies that outperform baselines on realistic image ordering tasks.
-Both theoretical and empirical results highlight the benefit of accounting for aleatoric uncertainty when learning from human annotators.
-The optimality of our sampling strategies remains an open question.
+We derived an upper bound on the ordering error and used it to design an active sampling strategy that outperforms or matches baselines on realistic image and text ordering tasks.
+Both theoretical and empirical results highlight the benefit of accounting for noise in comparisons when learning from human annotators.
+The optimality of our sampling strategy remains an open question.
 A future direction is to derive a lower bound on the ordering error, and prove an—ideally matching—algorithm-specific upper bound.
-Furthermore, for now, the benefit of utilizing item attributes depends on the quality of features that can be extracted using a pre-trained model.
-This could most likely be improved by performing representation learning throughout the learning process.
-Finally, while we perform experiments with real-world human annotations, these are limited to already-collected (offline) data.
-Future work should evaluate the strategies in an online setting where participants annotate comparisons recommended by the algorithms.
+However, constructing upper bounds for related fixed-budget tasks is an open problem (Qin, 2022).
+Moreover, motivated by the annotation setting, our focus has been on reducing sample complexity and we leave it to future work to explore potential linear approximations of the sampling criteria and other trade-offs between sample complexity and computational complexity.
+Further, our approach can potentially be improved by performing representation learning throughout the learning process.
+Finally, our experiments are constrained to a limited amount of already-collected (offline) human preference data, causing different algorithms to select disproportionately similar comparisons.
+Future work should evaluate the strategies in an online setting.
+Acknowledgements
+FDJ and HB are supported by Swedish Research Council Grant 2022-04748.
+FDJ is also supported in part by the Wallenberg AI, Autonomous Systems and Software Program founded by the Knut and ALice Wallenberg Foundation.
+EC and DD are supported by Chalmers AI Research Centre (CHAIR).
+The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no.
+2022-06725.
 Appendix A Notation
@@
 $f$ Comparison logit (typically linear), e.g., $f_{\theta}(i,j)=\theta^{\top}(x_{i}-x_{j})$
-Appendix B Extended related work
-Preference learning aims at ordering a set of items based on (noisy) outcomes of pairwise comparisons (Fürnkranz and Hüllermeier, 2003; Chu and Ghahramani, 2005).
-The problem is related also to learning to rank  (Burges et al., 2005; Jamieson and Nowak, 2011; Busse et al., 2012).
-Adaptively collecting data to solve preference learning with as few examples as possible may be posed as an active learning (Brinker, 2004; Silva et al., 2014; Long et al., 2010; Ling et al., 2020) or bandit problem.
-An example of the former, Houlsby et al.
-(2011) proposed the algorithm and applied it to active preference learning with per-item context variables and noisy pairwise preference feedback, adaptively proposing items for comparison by maximizing the expected information gain which is a commonly used acquisition function in Bayesian optimal design (Lindley, 1956).
-Singh and Chakraborty (2021) study age estimation through active learning of deep neural networks, like our task in Section 6.2, but used a mix of absolute and relative feedback.
-Bandit algorithms with relative feedback are often referred to as dueling bandits (Yue and Joachims, 2009; Bengs et al., 2021; Yan et al., 2022) and have been proposed both for items with context attributes (Dudík et al., 2015) and without (Yue et al., 2012).
-Recently, Bengs et al.
-(2022) proposed CoLSTIM which learns from relative feedback in the linear stochastic transitivity model, matching the model in (2) and Di et al.
-(2023) gave variance-aware regret bounds for stochastic contextual dueling bandits.
-The bandit paradigm closest to our problem is fixed-budget pure-exploration (Fang, 2022), where the goal is to identify the highest-ranked item in a set using a fixed budget of comparisons.
-Related to our work is also George and Dimitrakakis (2023) who consider the problem of finding a Kemeny ranking in non-contextual dueling bandits.
-Another variant is minimization of Borda regret (Wu et al., 2023b).
-Jun et al.
-(2021) considers a logistic model with pairwise comparisons where the goal is to find the item with highest ranking and they provide guarantees in the fixed-confidence regime.
-However, as we argue in Appendix E, a model that identifies the highest-ranked item optimally fast can be arbitrarily bad at recovering a total ordering.
-Our work is also related to the generalized linear bandit (Filippi et al., 2010; Li et al., 2017; Faury et al., 2020; Kveton et al., 2020) and our theoretical analysis leverage techniques from this literature.
-Another view of the preference learning problem is offered by non-contextual ranking methods, such as TrueSkill (Herbrich et al., 2006; Minka et al., 2018), Hamming-LUCB (Heckel et al., 2018) and Probe-Rank (Lou et al., 2022).
-These produce an in-sample ordering of a fixed set of items but must be updated if new items are added.
-We are also interested in the theoretical analysis of active preference learning.
-Chaudhuri et al.
-(2015) give an algorithm for and analysis of active learning for general maximum likelihood estimation.
-Ash et al.
-(2021) extend this idea and propose a tractable algorithm for models parameterized by deep neural network.
-Kirsch and Gal (2022) show that many recently proposed active learning strategies may be unified in a framework based on Fisher information.
-Preference learning is commonly used when training large language models through reinforcement learning with human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Wu et al., 2023a).
-In this line of work, Zhu et al.
-(2023) is closest to our setting and they provide guarantees on the sample complexity of learning a preference model from offline data.
-They leverage similar tools from statistical learning and bandits as we do.
-In contrast to their work, we provide sampling strategies and guarantees for the online setting.
-Mehta et al.
-(2023) consider active learning for RLHF in a dueling bandit framework where the goal is to optimize a contextual version of the Borda regret.
-Appendix C Algorithms
-C.1 MLE estimator for logistic regression
+Appendix B Algorithms
+B.1 MLE estimator for logistic regression
 The log-likelihood $L_{t}(\theta)$ of data $D_{t}=\{(i_{s},j_{s},c_{s})\}_{s=1}^{t}$, with $z_{s}=x_{i_{s}}-x_{j_{s}}$, under a logistic regression model with parameters $\theta$ is defined by
@@
 $\theta^{R}_{t}=\operatorname*{arg\,min}_{\theta}-L_{t}(\theta)+\lambda\|\theta\|_{2}^{2}~{}.$
-C.2 Bayesian estimator for logistic regression
+B.2 Bayesian estimator for logistic regression
 $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ according to the log likelihood
@@
 The hessian at time $t$ is defined as
-$\mathbf{H}_{B,t}=\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta)z_{i,j}z_{i,j}^{\top}=\mathbf{H}_{B,0}+\mathbf{H}_{t}.$
-Note that if priors $\theta_{B,0}=\mathbf{0}$ and $\mathbf{H}^{-1}_{B,0}=I_{d}$ are used, the log likelihood boils down to:
+$\mathbf{H}_{B,t}=\mathbf{H}_{B,0}+\sum_{(i,j)\in D_{t}}\dot{\sigma}(z_{i,j}^{\top}\theta_{B,t})z_{i,j}z_{i,j}^{\top}=\mathbf{H}_{B,0}+\mathbf{H}_{t}.$
+Moreover, if priors $\theta_{B,0}=\mathbf{0}$ and $\mathbf{H}^{-1}_{B,0}=I_{d}$ are used, the log likelihood boils down to:
 $\begin{split}\ln p(\theta\mid D_{t})=&-\frac{1}{2}||\theta||_{2}^{2}+\sum_{t}c_{t}\ln(\sigma(z_{i_{t},j_{t}}^{\top}\theta))+(1-c_{t})\ln(1-\sigma(z_{i_{t},j_{t}}^{\top}\theta))+const\end{split}$
-which implies that the MAP estimate will be the same as the MLE estimate in the frequentist setting with a ridge regularisation. Similarly the Hessian becomes:
+which implies that the MAP estimate will be the same as the MLE estimate with ridge regularisation in the frequentist setting. Similarly, the Hessian becomes:
 $\mathbf{H}_{B,t}=\mathbf{H}_{t}+I_{d}$
-C.3 Uniform sampling
+Sequential updates are also possible in the Bayesian setting by using your current estimates as the new priors.
+Note that this will give slightly different results, as the calculation of $\mathbf{H}_{B,t}$ depends on the current estimate of $\theta_{B,t}$.
+B.3 Stochastic Bayesian uncertainty reduction (BayesGURO)
+We describe BayesGURO, a Bayesian sampling criterion, closely related to GURO.
+Consider a Bayesian model of the parameter $\theta$ with $p(\theta)$ the prior belief and $p(\theta\mid D_{t})$ the posterior after observing the preference feedback in $D_{t}$.
+A natural strategy for learning more about the ordering of $\mathcal{I}$ is to sample items $i_{t},j_{t}$ based on an estimate of the posterior variance of predictions for their comparison,
+$i_{t},j_{t}=\operatorname*{arg\,max}_{i,j\in\mathcal{I}_{D},i<j}\hat{\mathbb{V}}_{\theta\mid D_{t-1}}[\sigma(\theta^{\top}z_{ij})]~{}.$ (12)
+Here, $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ is an estimate of the variance of probabilities $\sigma(\theta^{T}z_{ij})$, computed from finite samples drawn from the posterior of $\theta$.
+Estimating the variance in this way both i) allows for tractable implementation, and ii) induces controlled stochasticity in the selection of item pairs.
+This can be useful in batched learning settings so that multiple pairs can be sampled within the same batch.
+A deterministic criterion would return the same item pair every time until $\theta$ is updated.
+We refer to the sampling criterion in (7) as BayesGURO.
+For the logistic model considered in Section 4, using Laplace approximation with a Normal prior $\mathcal{N}(0,\mathbf{H}^{-1}_{B,0})$ on $\theta$, the Bayesian criterion in (7) is related to the GURO sampling criterion in (5) through the first-order Taylor expansion of the variance:
+$\mathbb{V}_{\theta\mid D_{t}}(\sigma(\theta^{\top}z_{ij}))$ $\approx(\dot{\sigma}(\mathbb{E}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]))^{2}\mathbb{V}_{\theta\mid D_{t}}[\theta^{\top}z_{ij}]=(\dot{\sigma}(\theta_{B,t}^{\top}z_{ij})||z_{ij}||_{\mathbf{H}^{-1}_{B,t}(\theta_{B,t})})^{2}~{},$
+where $\theta_{B,t}$ is the MAP estimate of $\theta$ at time $t$ and $\mathbf{H}_{B,t}$ is the Hessian adjusted by the prior covariance $\mathbf{H}^{-1}_{B,0}$ (further described in Appendix B.2).
+Thus, to a first-order approximation, for a large number of posterior samples, the GURO and BayesGURO active learning criteria are equivalent, save for the influence of the prior.
+In practice, we find that the Bayesian variant lends itself well to sequential updates of the posterior.
+The choice of prior $p(\theta)$, which could be useful under strong domain knowledge, and the stochasticity of using few posterior samples to approximate $\mathbb{V}$ make the two criteria distinct.
+B.4 Uniform sampling
 The uniform sampling algorithm is given in Algorithm 2. The corresponding Bayesian version replaces line 5 with the MAP estimate.
@@
 8:Return $h_{T}$
-C.4 BALD
+B.5 BALD
 Algorithm 3 BALD bandit
@@
 10:Return $h_{T}$
-Where the posterior is calculated as in Appendix C.2 and $H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ is approximated as in Appendix C.4.1.
-C.4.1 Deriving the BALD sampling criterion
+Where the posterior is calculated as in Appendix B.2 and $H[y\mid z_{i,j},D_{t-1}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t-1})}[H[y\mid z_{i,j},\mathbf{\theta}]]$ is approximated as in Appendix B.5.1.
+B.5.1 Deriving the BALD sampling criterion
 The BALD criteria formalized using our notation becomes
@@
 Here $\Pr(y\mid z_{i,j},D_{t})$ is the predictive distribution for our Bayesian logistic regression model.
-As covered in Bishop and Nasrabadi (2006), this expectation cannot be evaluated analytically but can be approximated using the probit function $\Phi$;
+As covered in Bishop and Nasrabadi (2006, Chapter 4), this expectation cannot be evaluated analytically but can be approximated using the probit function $\Phi$;
 $\Pr(y\mid z_{ij},D_{t})\approx\Phi\left(\frac{\mathbf{\theta}^{\top}_{t}z_{i,j}}{\sqrt{\lambda^{-2}+||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\approx\sigma\left(\frac{\mathbf{\theta}^{\top}_{t}z_{i,j}}{\sqrt{1+\frac{\pi||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}(\theta_{*})}}{8}}}\right).$
@@
 […]
-where $C=\sqrt{4\ln{2}}$. Finally, we arrive at an estimation of the objective function which we wish to maximize:
-$H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]\approx h\left(\sigma\left(\frac{\theta_{t}^{\top}z_{i,j}}{\sqrt{1+\frac{\pi}{8}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\right)-\frac{C}{\sqrt{||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2}}}\exp{\left(-\frac{(\theta_{t}^{\top}z_{i,j})^{2}}{(||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2})}\right)}$
-Appendix D Proofs of Lemma 1 and Theorem 1
-D.1 Proof of Lemma 1
+where $C=\sqrt{4\ln{2}}$. Finally, we arrive at an estimation of the objective function we wish to maximize:
+$H[y\mid z_{i,j},D_{t}]-\mathbb{E}_{\mathbf{\theta}\sim p(\mathbf{\theta}\mid D_{t})}[H[y\mid z_{i,j},\mathbf{\theta}]]\approx$ $\ h\left(\sigma\left(\frac{\theta_{t}^{\top}z_{i,j}}{\sqrt{1+\frac{\pi}{8}||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}}}\right)\right)$
+$-\frac{C}{\sqrt{||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2}}}\exp{\left(-\frac{(\theta_{t}^{\top}z_{i,j})^{2}}{(||z_{ij}||^{2}_{\mathbf{H}^{-1}_{t}}+C^{2})}\right)}$
+Appendix C Proofs of Lemma 1 and Theorem 1
+C.1 Proof of Lemma 1
 Proof.
@@
 Using the definition of $\tilde{\mathbf{H}}_{t}$ yields the statement of Lemma 1. ∎
-D.2 Proof of Theorem 1
+C.2 Proof of Theorem 1
 Proof.
-We let $i\succ j$ denote that $i$ is preferred to $j$.
-W.l.o.g assume $1\succ 2\succ...\succ n$.
-The key observation is that the assumption $\sigma(z_{ij}^{\top}\theta_{*})-0.5>\Delta_{ij}>\Delta_{*},\forall i,j$ implies that for any $i$ and $j$ such that $i<j$ it holds that
+We let $i\succ j$ denote that $i$ is preferred to $j$. W.l.o.g assume $1\succ 2\succ...\succ n$. The key observation is that for any $i$ and $j$ such that $i<j$ it holds that
 $\Delta_{i,j}>(j-i)\Delta_{*}.$
@@
 $P(R(\theta_{T})\geq\epsilon)\leq\frac{4dT}{\epsilon n}\left(\left(\alpha_{*}^{-1}-1\right)^{-1}+\left(\beta_{*}^{-1}-1\right)^{-1}\right).$ (25)
-Appendix E Comparison with regret minimization
+C.3 Extensions of current theory
+Regularized estimators.
+In our analysis in Section 4, we have assumed that $\theta_{T}$ is the maximum likelihood estimate and that $\mathbf{H}(\theta_{T})$ has full rank.
+This can be relaxed by considering $\ell_{2}$ (Ridge) regularization where $\theta_{\lambda_{0},T}$ is the optimum of the regularized log-likelihood with regularization $\lambda_{0}\mathbf{I}$ and $\mathbf{H}_{\lambda_{0}}(\theta_{\lambda_{0},T})=\sum_{s=1}^{T}\dot{\sigma}(z_{s}^{\top}\theta_{\lambda_{0},T})z_{s}z_{s}^{\top}+\lambda_{0}\mathbf{I}$.
+The same machinery used to prove Lemma 1 (Filippi et al., 2010; Faury et al., 2020) can be applied to this regularized version with small changes to the final bound.
+Generalized linear models.
+It is also possible to derive similar results for generalized linear models with other link functions, $\mu(z_{ij}^{\top}\theta_{*})$, by using the general inequality $\mathbf{H}(\theta)\geq\kappa^{-1}\mathbf{V}$ with $\mathbf{V}=\sum_{s=1}^{T}z_{s}z_{s}^{\top}$ and $\kappa\geq 1/\min_{z_{ij}}\dot{\mu}(z_{ij}^{\top}\theta_{*})$.
+We conjecture that this will yield a scaling of $\sim\exp(-\Delta^{2}T/\kappa)$ where, unfortunately, $\kappa$ might be very large.
+For a more thorough discussion on the dependence on $\kappa$ in generalized linear bandits, see Lattimore and Szepesvári (2020, Chapter 19).
+Lower and algorithm-specific upper bounds on the ordering error.
+A worst-case lower bound on the ordering error can be constructed in the fixed-confidence setting, where the goal is to minimize the number of comparisons until a correct ordering is found with a given confidence, by following Garivier and Kaufmann (2016).
+This involves defining the set of alternative models $\mathrm{Alt}(\theta_{*})$ which differs from $\theta_{*}$ in their induced ordering of $\mathcal{I}$.
+The bound is then constructed by optimizing the frequency of comparisons of each pair of items so that such alternative models are distinguished as much as possible from the true parameter.
+We have left this result out of the paper as we find it uninformative in the regime when the number of comparisons is small, (see Simchowitz et al.
+(2017) for a discussion on the limitations of these asymptotic results in the standard bandit setting).
+Constructing a lower bound for our fixed-budget setting, of learning as good an ordering as possible with a fixed number of comparisons, is much more challenging.
+The fixed-confidence result yields a bound for the fixed-budget case (Garivier and Kaufmann, 2016), but constructing either a tight lower bound or a tight algorithm-specific upper bound is an open problem (Fang, 2022).
+Appendix D Comparison with regret minimization
 Bengs et al.
@@
 Saha and Krishnamurthy (2022) investigated where there is a computationally efficient algorithm that achieves the derived optimality guarantee.
-Appendix F Experiment details
-F.1 Ground truth on PhotoAge
-While the IMDB dataset used in the PhotoAge task contains an associated age with every image, these were not used.
-There were two reasons for this.
-First, some of these ages were incorrectly assigned (often in photos containing multiple people that had been cropped to one individual), causing some noise.
-Second, and most importantly, we are interested in the subjective assessment of how old someone appears to be, which does not have to be consistent with the true age.
-The consequence of this is that if you keep collecting comparisons, there is no guarantee that the ordering will ever converge to reflect the true age.
-The rationale behind using the Borda Score instead is that the simple algorithm does not make any assumptions regarding transitivity Nihar B.
-Shah (2017), thus not favoring any of the sampling strategies.
-F.2 Implementation Details PhotoAge
-To increase computational efficiency, the hybrid algorithms did not evaluate all 100 000similar-toabsent100000$\sim$ 100\ 000∼ 100 000 comparisons at every time step.
-Instead, a subset of 5 00050005\ 0005 000 comparisons was first sampled, and the highest scoring pair in this set was chosen.
-This resulted in a large speed up, and no noticeable change in performance during evaluation.
-F.3 Additional figures
-NormMin Comparison
+Appendix E Experiment details
+For BayesGURO and BALD, the posterior $p(\theta\mid D_{t})$ is estimated using the Laplace approximation as described in Bishop and Nasrabadi (2006, Chapter 4).
+With this approximation, the covariance matrix is the same as the inverse of the Hessian of the log-likelihood.
+For both methods, the priors $\theta_{B,0}=\mathbf{0}^{d}$ and $\mathbf{H}_{B,0}^{-1}=I_{d}$ were used, and sequential updates were performed every iteration.
+The sample criterion for BALD under a logistic model is given in Appendix B.5.1.
+For BayesGURO, 50 posterior samples were used to estimate $\hat{\mathbb{V}}_{\theta\mid D_{t}}[\sigma(\theta^{T}z_{ij})]$ for every $z_{ij}$.
+The hybrid algorithms follow the same structure with the added constraint that each per-item parameter $\zeta_{i}$ is independent of other parameters.
+This allows for efficient updates of $\mathbf{H}_{B,t}^{-1}$ by using sparsity in the covariance.
+GURO, CoLSTIM, and Uniform use LogisticRegression from Scikit-learn (Pedregosa et al., 2011) with default Ridge regularization ($C=1$) and the lbfgs optimizer.
+The former two updates $\theta_{t}$ every iteration using the full history, $D_{t}$ in all experiments except for IMDB-WIKI-SbS, where GURO updates $\theta_{t}$ every 25th iteration.
+This caused no noticeable change in performance as GURO still updates $\mathbf{H}^{-1}_{t}$ every iteration using the Sherman-Morrison formula.
+Note that when using the Sherman-Morrison formula in practice, you only get an estimate of $\mathbf{H}^{-1}_{t}(\theta_{t})$ since previous versions have been calculated using older estimates of $\theta$.
+This method for approximating the inverse hessian is covered in Bishop and Nasrabadi (2006, Chapter 5) and when we compared it to calculating $\mathbf{H}^{-1}_{t}(\theta_{t})$ from scratch every iteration we observed that the methods performed equally.
+The design matrix for CoLSTIM is updated as in Bengs et al.
+(2022): the confidence width $c_{1}$ was chosen to be $\sqrt{d\log(T)}$, and the perturbed values were generated using the standard Gumbel distribution.
+To increase computational efficiency for the large IMDB-WIKI-SbS dataset, the hybrid algorithms did not evaluate all 100 000similar-toabsent100000$\sim$ 100\ 000∼ 100 000 comparisons at every time step.
+Instead, a subset of 5 00050005\ 0005 000 comparisons was first sampled, and the highest-scoring pair in this set was chosen.
+This resulted in a large speed-up and no noticeable change in performance during evaluation.
+E.1 Datasets
+ImageClarity
+Data available at https://dbgroup.cs.tsinghua.edu.cn/ligl/crowdtopk.
+This dataset contained differently distorted versions of the same image.
+To extract relevant features, we used a ResNet34 model (He et al., 2016) that had been pre-trained on Imagenet (Deng et al., 2009).
+After PCA projection feature dimensionality was reduced to $d=63$.
+The dataset consisted of $100$ images and 27 7302773027\ 73027 730 comparisons.
+Since the type of distortion is the same for all images, the dataset has a true ordering with regards to the strength of the distortion applied.
+WiscAdds
+Data available at https://dataverse.harvard.edu/dataset.xhtml?persistentId= doi:10.7910/DVN/0ZRGEE (license: CC0 1.0).
+The WiscAdds dataset, containing $935$ political texts, has been extended with 9 52895289\ 5289 528 pairwise comparisons by Carlson and Montgomery (2017).
+In comparisons, annotators have stated which of two texts has a more negative tone toward a political opponent.
+To extract general features from the text, sentences were embedded using the pre-trained all-mpnet-base-v2 model from the Sentence-Transformers library (Reimers and Gurevych, 2019).
+After applying PCA to the sentence embeddings, each embedding had a dimensionality of $d=162$.
+IMDB-WIKI-SbS
+Data available at https://github.com/Toloka/IMDB-WIKI-SbS (license: CC BY).
+IMDB-WIKI-SbS consists of close-up images of actors of different ages.
+For each comparison, the label corresponds to which of two people appears older.
+The complete dataset consists of 9 15091509\ 1509 150 images and 250 249250249250\ 249250 249 comparisons, but images that were grayscale or had a resolution lower than $160\times 160$ were removed, resulting in 6 072 images and 110 349 comparisons.
+We extract features from each image using the Inception-ResNet implemented in FaceNet (Schroff et al., 2015) followed by PCA, resulting in $d=75$ features per image.
+E.2 Additional figures
+X-RayAge
 To highlight the importance of the first-order term in Lemma 1, we evaluated NormMin on the same X-ray ordering task as in Figure 1(a).
-The results, shown in Figure 5, indicate that not only does the algorithm perform worse than GURO, but is seemingly also outperformed by a uniform sampling strategy.
-Figure 5: X-RayAge. NormMin included in the experiment shown in Figure. 1(a)
+The results, shown in Figure 7(a), indicate that not only does the algorithm perform worse than GURO, but is seemingly also outperformed by a uniform sampling strategy.
+Furthermore, for completeness, we include Figure 7(b) which shows the in-sample error, $R_{I_{D}}$, during the generalization experiment.
+(a) X-RayAge. NormMin included in the experiment shown in Figure 1(a).
+(b) X-RayAge. The in-sample error $R_{I_{D}}$ for the generalization experiment performed in Figure 1(b).
+Figure 7: Additional figures from the X-RayAge experiment.
 Synthetic Example and Illustration of Upper Bound
@@
 Randomly initialized representation
-As discussed in Section 6.2, the performance of our contextual approach will depend on the quality of the representations.
-To underscore the practical usefulness of our algorithms, we have performed the same experiment as in Figure 4, but this time the model used to extract image features was untrained (i.e.
-the weights were random).
+As discussed in Section 1, the performance of our contextual approach will depend on the quality of the representations.
+To underscore the practical usefulness of our algorithms, we have performed the same experiment as in Figure 6, but this time the model used to extract image features was untrained (i.e., the weights were random).
 As to be expected, the results, shown in Figure 9, demonstrate that the fully contextual algorithms have no real way of ordering the items according to these uninformative features.
-However, the hybrid BayesGURO algorithm performs similarly to TrueSkill, despite model misspecification.
+However, GURO Hybrid performs similarly to TrueSkill, despite model misspecification.
 This is promising, since you may not know in advance how informative the extracted features will be for the target ordering task.
-Figure 7: PhotoAge. The same experiment as presented in Figure. 4, but the model used for feature extraction is untrained.
+Figure 9: IMDB-WIKI-SbS. The same experiment as presented in Figure. 6, but the model used for feature extraction is untrained.
+ImageClarity ground truth
+The ImageClarity dataset consists of multiple versions of the same image, with the same distortion applied to it to varying degrees.
+Due to this artificial construction, the pairwise comparisons should, given enough samples, reflect the magnitudes of the applied distortions.
+In Figure 10 we perform the same experiment as in Figure 6, but instead of evaluating on a holdout comparison set, we measure the distance to the ground-truth ordering.
+The overall results are very similar, although we do see a slight increase in the performance of contextual algorithms compared to the non-contextual TrueSkill.
+Figure 10: ImageClarity. Same experiment as in Figure 6, but now measuring the distance to the ground-truth ordering. Averaged over 25 seeds along with the 1-sigma error region.
+Ground truth ordering using the Bradley-Terry model
+An alternate approach to evaluate ordering quality is to estimate a ”ground-truth” ordering by applying the popular Bradley-Terry (BT) model (Bradley and Terry, 1952) to all available comparisons.
+We used the CrowdKit library (Ustalov et al., 2024) to find the MLE scores for each item and ordered the elements accordingly.
+In Figure 14 we run the same experiments as in Figure 6, but instead measure the distance to the constructed BT ordering.
+The overall trends remain, but for (14) and (14) there is a slight shift for the later iterations.
+More specifically we see non-contextual TrueSkill eventually overtaking the contextual algorithms.
+{subcaptionblock}
+[t]0.49 Refer to caption {subcaptionblock}[t]0.49 Refer to caption
+Figure 11: ImageClarity. $n=100$, $d=63$.
+Figure 12: WiscAds. $n=935$, $d=162$.
+{subcaptionblock}
+[t]0.49 Refer to caption
+[…]
+Figure 14: The same experiment as presented in Figure. 6, but we instead measure the distance to a ground-truth estimated using all available comparisons.
+The issue is that algorithms with orderings closer to the maximum likelihood estimate of the BT model will be favored.
+To exemplify this we use the ImageClarity dataset since it contains the largest number of comparisons relative to the number of items.
+We sample 1 00010001\ 0001 000 comparisons and let this be the collection that is available to the algorithms.
+We further construct two target orderings, one from the BT estimate using the sampled subset of comparisons, and a second, more probable ordering, from the BT estimate using all 27 7302773027\ 73027 730 available comparisons.
+Figure 15 shows the distance between the GURO and TS algorithms and the different target orderings, where dashed lines indicate the distance to the ordering generated using the full list of comparisons.
+If we only look at the distance to the ordering produced using our subset of comparisons, TrueSkill seemingly outperforms GURO after about $350$ comparisons.
+However, if we instead measure the distance to the more probable ordering, we see that GURO converges toward a lower distance.
+Note that these are the same orderings, evaluated against different targets.
+This is likely the effect we observe in Figure 14 and 14, but not in Figure 14 as a result of the high amount of comparisons available to us.
+Figure 15: ImageClarity.
+The same experiment as presented in Figure 6, but we instead measure the distance to target orderings that correspond to the maximum likelihood estimate of the BT model using different numbers of comparisons.
+The dashed lines show the distance to the BT estimate using all 27 7302773027\ 73027 730 comparisons.